### v2

# Carbon-Efficient Software Design and Development: A Systematic Literature Review

Abstract.
The ICT sector, responsible for 2% of global carbon emissions, is under scrutiny calling for methodologies and tools to design and develop software in an environmentally sustainable-by-design manner. However, the software engineering solutions for designing and developing carbon-efficient software are currently scattered over multiple different pieces of literature, which makes it difficult to consult the body of knowledge on the topic. In this article, we precisely conduct a systematic literature review on state-of-the-art proposals for designing and developing carbon-efficient software. We identify and analyse 65 primary studies by classifying them through a taxonomy aimed at answering the 5W1H questions of carbon-efficient software design and development. We first provide a reasoned overview and discussion of the existing guidelines, reference models, measurement solutions and techniques for measuring, reducing, or minimising the carbon footprint of software. Ultimately, we identify open challenges and research gaps, offering insights for future work in this field.
1. Introduction
In recent years, the themes of sustainability and sustainable development have seen a growing interest from governments, academia and industry, aiming at balancing the need to safeguard our planet along with social and economical instances (Brundtland et al., 1987; Sachs et al., 2022). Such effort is clearly epitomised by the adoption of the Sustainable Development Goals in the UN action plan reported in the 2030 Agenda for Sustainable Development.111Publicly available at: https://sdgs.un.org/2030agenda Indeed, global carbon emissions need to be reduced to secure our Planet’s future by achieving the ambition set by the Paris agreement222https://www.un.org/en/climatechange/paris-agreement (Manner, 2022; Möller et al., 2024). However, such an objective remains far from being reached, as highlighted by the outcomes of the 2023 United Nations Climate Change Conference (COP28).
Building on this momentum, there is a rising concern about measuring, assessing, and improving the sustainability of Information and Communication Technology (ICT). Indeed, ICT is currently responsible for 2% of global carbon emissions (Manner, 2022; Mingay and Pamlin, 2008), with an energy consumption estimated at between 6% and 9% of the global demand and expected to grow up to 20% by 2030 (Belkhir and Elmeligi, 2018; Zulfiqar et al., 2023; Adisa et al., 2024). Most efforts to reduce the carbon footprint of ICT were dedicated to improve hardware energy-efficiency, by leveraging the combined effects of Moore’s law (Moore, 1998) and Dennard scaling (Dennard et al., 1974). Moore’s law ensured the possibility of packing double the transistors in a chip, while Dennard scaling has ensured that integrated circuits of the same area maintain constant power consumption regardless of the increased number of transistors. This resulted in faster chips with no increase in energy consumption up until the 2010s (Anderson et al., 2023).
With Dennard’s scaling and Moore’s law reaching their limits (Muralidhar et al., 2022), improving the environmental sustainability of ICT will necessarily involve making software more sustainable by reducing the carbon footprint associated with its deployment. On this line, much research has targeted the optimisation of renewable energy usage through suitable service orchestration and consolidation (Gaglianese et al., 2023a), without considering the potential of (re)designing and (re)writing software according to a sustainable-by-design approach. Naturally, as pointed out by Anderson et al. (Anderson et al., 2023), further improvements in software sustainability will require treating carbon-efficiency in our code design and development. Similarly, Verdecchia et al. (Verdecchia et al., 2021) call for methodologies to write sustainable software by blending decision-making with the software execution context. Although the term green software engineering was coined more than a decade ago (Caballar, 2024), much remains to be done to fully rethink how carbon-efficient software is designed and implemented.
At the same time, as observed, e.g., by the Green Software Foundation (Murugesan, 2021) and Kern et al. (Kern et al., 2015), the carbon-efficiency of software products is mostly affected by their design, development, and operation. Indeed, while sustainability considerations span the entire software lifecycle, design, development, and operation represent the primary leverage points for reducing software carbon emissions, e.g., through software architecture decisions, efficient data structures and algorithms, and energy- and carbon-aware orchestration strategies.
While the environmentally sustainable operation of software products has already been analysed, e.g., by Gaglianese et al. (Gaglianese et al., 2023b), this article precisely aims at surveying the state of the art related to the design and development of carbon-efficient software. Currently available methodologies are scattered over multiple different pieces of literature. Unfortunately, this makes it difficult to consult the body of knowledge on the topic, both for researchers willing to investigate carbon-efficient software design and development methodologies, and for practitioners willing to apply them to their software systems. To achieve this objective, this article conducts a systematic literature review (SLR), which identifies, taxonomically classifies, and compares the existing guidelines, reference models, measurement solutions, and techniques for designing and developing carbon-efficient software. Following the guidelines by Petersen et al. (Petersen et al., 2008, 2015) on how to conduct SLRs in software engineering, we first identify a corpus of 65 primary studies. We then classify and compare the selected studies by means of a newly proposed taxonomy, aimed at highlighting the 5W1H of carbon-efficient software design and development, viz., who are the stakeholders targeted by the selected studies, what methods are proposed therein, why, where and when the proposed methods are applied (in terms of application domain and software lifecycle step, respectively), and how the selected studies have been released to the community.
Overall, our SLR pursues a twofold ultimate objective. On one hand, it aims at accurately depicting existing proposals, methodologies, and solutions that enable the design and development of environmentally sustainable software, namely for measuring, reducing, and/or minimising a software’s carbon emissions. On the other hand, it analyses the state of the art in carbon-efficient software design and development with the aim of shedding light on open challenges and gaps, which can contribute to reducing the carbon footprint of ICT from a software perspective.
The rest of this article is organised as follows. Section 2 describes the design of the SLR process. Section 3 classifies and summarises the state of the art on carbon-efficient software design and development. The classification is retaken in Section 4, which provides an analysis of the 5W1H of designing and developing carbon-efficient software. Sections 5 and 6 discuss the possible threats to the validity of our SLR and related work, respectively. Finally, Section 7 concludes the article by highlighting future research directions for the design and development of carbon-efficient software.
2. Research Design
After explicitly stating the objective and the 5W1H questions that drove our SLR (Section 2.1), we illustrate the design of the SLR (Sections 2.2, 2.3 and 2.4) by relying on Figure 1 to provide a graphical sketch of the SLR process.
We also provide a replication package for repeating our SLR (Section 2.5).
2.1. The 5W1H Questions
The objective of this article is to identify, classify, and analyse the state of the art in the design and development of environmentally sustainable software by limiting its carbon footprint. This is done by answering to the 5W1H questions listed below:
Who? Who are the stakeholders in existing techniques for carbon-efficient software design and development?
What? What methods, tools, and techniques have been proposed to achieve carbon-efficient software design and development?
Why? Why does a selected study target the carbon efficiency of software? Does it aim at measuring, reducing, or optimising its carbon footprint?
Where? In which application domains can a selected study be applied?
When? In which software lifecycle stages can a selected study’s proposal be used?
How? How were the selected studies published?
2.2. Literature Search
Following the guidelines by Petersen et al. (Petersen et al., 2008, 2015), we identified the search string guided by our main research objective, namely to overview the state of the art in carbon-efficient software design and development. More precisely, we defined the search strings based on the PICO (Population, Intervention, Comparison, Outcome) terms characterising our research questions (Kitchenham and Charters, 2007), taking keywords from each aspect of our research objective – and the corresponding 5W1H questions. Differently from Petersen et al. (Petersen et al., 2008, 2015), we did not limit our focus to specific outcomes or experimental designs in our study. Rather, we aimed at providing a broader overview of the state of the art in the design and development of carbon-efficient software. This difference is also reflected in the search string we employed to search for studies, i.e.,
(where “*” matches lexically related terms, e.g., plurals and conjugations, and relevant synonyms). The search string is essentially the conjunction of three groups, aimed at identifying the existing efforts (1) considering carbon emissions, (2) reducing or containing such carbon emissions, and (3) focusing on software design and development.
We used the aforementioned string to search for all white literature published before January 2024, when this research began. The search was carried on ACM Digital Library, IEEEXplore, Scopus, and Web of Science, and was restricted to white literature published in Computer Science venues. We retrieved a total of 1517 candidate studies, which – after checking for duplicates – were reduced to a list of 1273 unique candidate studies.
2.3. Literature Selection
The search terms used in Section 2.2 were checked against the titles and abstracts of the existing literature, which resulted in a significant number of candidate studies that were not relevant for our purposes. Therefore, the list of candidate studies was further refined by relying on the inclusion/exclusion criteria in Table 1. The inclusion criteria and aligned the selection with our main research objective by allowing the selection of existing studies that support the design and/or development of software, while also aiming at reducing its carbon footprint. The exclusion criteria , , and restricted the focus to peer-reviewed literature written in English, which was also accessible from our institution. Finally, the exclusion criteria and restricted the focus to primary studies and excluding collections of papers and secondary studies, respectively. A candidate study was selected if satisfying all the inclusion criteria, whereas it was excluded if satisfying any of the exclusion criteria.
[…]
| Inclusion | The study explicitly targets reducing the carbon footprint of software | |
[…]
| Exclusion | The study is not peer-reviewed | |
[…]
| Exclusion | The study is not accessible from our institution | |
[…]
| Exclusion | The study is a secondary study (e.g., a survey or SLR) |
The inclusion/exclusion criteria were applied in two consecutive steps. In both steps, each candidate study was marked as to be selected or excluded by two authors independently. If there was disagreement about the inclusion of a candidate study, the author who was not involved in the coding of the study acted as a third evaluator to resolve the disagreement. In the first step, the selection was based on assessing the inclusion/exclusion criteria over the title and abstract of candidate studies, which reduced the list of candidate studies to 88 candidate studies, with an overall inter-rater agreement of 85.6%.333We measured the inter-rater agreement on the inclusion/exclusion coding by exploiting the Krippendorff- coefficient (Krippendorff, 2004), which classifies a coding as good if the value is greater or equal to 80%. The process was then repeated by considering the full text of the remaining 88 candidate studies, which further narrowed the list to 43 candidate studies, with an overall inter-rater agreement of 90.9%.
[…]
| (Ahmed, 2020) Ahmed. Environmental sustainability coding techniques for cloud computing (2020) |
[…]
| (Alofi et al., 2022) Alofi et al. Optimizing the energy consumption of blockchain-based systems using evolutionary algorithms: a new problem … (2022) |
[…]
| (Becker et al., 2016) Becker et al. Requirements: the key to sustainability (2016) |
[…]
| (Brownlee et al., 2021) Brownleee et al. Exploring the accuracy-energy trade-off in machine learning (2021) |
[…]
| (Condori-Fernández and Lago, 2018) Condori-Fernàndez and Lago. The influence of green strategies design onyo quality requirement prioritization (2018) |
[…]
| (Dick et al., 2010) Dick et al. A model and selected instances of green and sustainable software (2010) |
[…]
| (Fu et al., 2021) Fu et al. Reconsidering CO2 emissions from computer vision (2021) |
[…]
| (Goaër, 2023) Le Goaër. Decarbonizinf software with free and open-source software: the ecoCode project (2023) |
[…]
| (Heithoff et al., 2023) Heithoff et al. Digital twins for sustainable software systems (2023) |
[…]
| (Ibrahim et al., 2021) Ibrahim et al. The development of green software process model: a qualitative design and pilot study (2021) |
[…]
| (Jayanthi et al., 2021) Jayanthi et al. An organizational structure for sustainable software development (2021) |
[…]
| (Johann et al., 2011) Johann et al. Sustainable development, sustainable software, and sustainable software engineering: an integrated approach (2011) |
[…]
| (Kern et al., 2013) Kern et al. Green software and green software engineering – definitions, measurements, and quality aspects (2013) |
[…]
| (Kern et al., 2018) Kern et al. Sustainable software products – Towards assessment criteria for resource and energy efficiency (2018) |
[…]
| (Kipp et al., 2011) Kipp et al. Green metrics for energy-aware IT systems (2011) |
[…]
| (Koçak and Alptekin, 2019) Kocak and Alptekin. A utility model for designing environmentally sustainable software (2019) |
[…]
| (Lago and Jansen, 2010) Lago and Jansen. Creating environmental awareness in service oriented software engineering (2011) |
[…]
| (Lami et al., 2014) Lami et al. An ISO/IEC 33000-compliant measurement framework for software process sustainability assessment (2014) |
[…]
| (Lasla et al., 2022) Lasla et al. Green-PoW: an energy-efficient blockchain proof-of-work consensus algorithm (2022) |
[…]
| (Mancebo et al., 2021) Mancebo et al. FEETINGS: Framework for energy efficiency testing to improve environmental goals of the software (2021) |
[…]
| (Mehra et al., 2022) Mehra et al. Towards a green quotient for software projects (2022) |
[…]
| (Moshnyaga, 2013a) Moshnyaga. An assessment of software lifecycle energy (2013) |
[…]
| (Moshnyaga, 2013c) Moshnyaga. Lifecycle energy assessment of mobile applications (2013) |
[…]
| (Naumann et al., 2015) Naumann et al. Sustainable software engineering: process and quality models, life cycle, and social aspects (2015) |
[…]
| (Pa et al., 2017) Pa et al. Dashboard system for measuring green software design (2021) |
[…]
| (Patterson et al., 2022) Patterson et al. The carbon footprint of machine learning will plateau, then shrink (2022) |
[…]
| (Ponsard et al., 2018) Ponsard et al. Building sustainable software for sustainable systems: case study of a shared pick-up and delivery service (2018) |
[…]
| (Raisian et al., 2021) Raisian et al. The green software measurement structure based on sustainability perspective (2021) |
[…]
| (Shenoy and Eeratta, 2011) Shenoy et al. Green software development model: an approach towards sustainable software development (2011) |
[…]
| (Sikand et al., 2023) Sikand et al. Green AI quotient: assessing greenness of AI-based software and the way forward (2023) |
[…]
| (Tee et al., 2014a) Tee et al. Toward developing green software development model in managing knowledge of IT practitioners for sustaining future… (2014) |
[…]
| (Wedyan et al., 2023) Wedyan et al. Integration and unit testing of software energy consumption (2023) |
[…]
| (Zhang et al., 2023) Zhang et al. Toward net-zero carbon emission in network AI for 6G and beyond (2023) |
The list of candidate studies was completed by running backward snowballing (Wohlin, 2014), which resulted in 30 additional candidate studies. These were again classified for inclusion/exclusion by repeating the same process as described above and by directly considering their full read. As a result, 22 additional studies were selected, reaching a total of 65 studies to be considered by our SLR. The final list of selected studies is shown in Table 2.
2.4. Literature Classification
The selected studies were classified by exploiting the taxonomy in Table 3. The taxonomy is rooted in the 5W1H questions driving our research (Section 2.1), with each of these questions being mapped to one or more possible classifications, i.e., categories of values. For the sake of brevity, Table 3 lists for each category the values associated with at least one selected study.
[…]
| Who? | stakeholder* | end user, IT operator, product owner, software architect, software provider, software developer, sustainability engineer | |
[…]
| Why? | goal* | indirectly | measure, reduce, optimise |
| directly | measure, reduce, optimise | ||
[…]
| Where? | target application domain | blockchain, cloud computing, computer vision, cyber-physical systems, edge computing, generic software, HPC, ML, mobile applications | |
[…]
Who? The taxonomy enables classifying the (potentially multiple) stakeholders for the selected studies. These include the end-user (who selects and uses a software product), the IT operator (who operates a software product), the product owner (who manages the software process and aims at maximising the value of a software product), the software architect (who designs of a software product), the software developer (who implements parts of a software product), the software provider (who delivers a software product to end-users), and the sustainability engineer (who engineers the sustainability of a software product).
What? The taxonomy in Table 3 enables classifying the proposal of each selected study by distinguishing among sustainability guidelines (in the form of shared good practices or experience reports), reference models (namely, abstract frameworks consisting of interlinked concepts aimed at clarifying what is carbon-efficiency for software design and development), measurement solutions (in the form of assistive tooling or estimate techniques supporting the measurement of a software’s sustainability), and software improvement techniques (namely, techniques aimed at supporting the design and/or development of software to reduce its carbon footprint).
The taxonomy also enables distinguishing whether the proposal of a selected study is white-box (or black-box), namely whether it requires the availabilty of a software’s source code, given its relevance to software design and development. It is also possible to indicate the TRL (Technology Readiness Level) of a selected study’s proposal from 1 to 9, following the TRL classification defined by the European Commission (European Commission, 2014).
Why? The taxonomy in Table 3 enables classifying the (potentially multiple) goals of the selected studies. Namely, it enables indicating whether a selected study aims at supporting the design and development of software while measuring, reducing, or optimising its carbon emissions. While reducing and optimising carbon emissions will certainly improve a software’s carbon efficiency, measurement can also support this, for example by comparing a software’s carbon footprint before and after updates. This is to check whether such updates have made the software more or less sustainable. We also distinguish whether this is done by directly measuring a software’s carbon emissions, or indirectly, by assuming that reducing a software’s energy consumption will also reduce its carbon footprint.
When? The taxonomy in Table 3 enables distinguishing the (potentially multiple) lifecycle stages supported by a selected study’s proposal. To this end, we follow the iterative software lifecycle process described by Kneuper (Kneuper, 2018), which consists of requirement analysis, design, implementation, integration & testing, and rollout, in that order.
Where? The taxonomy in Table 3 makes it possible to distinguish the target application domain for a selected study. These include blockchains, cloud computing, computer vision, cyber-physical systems, edge computing, HPC, ML, and mobile applications. The above listed domains are complemented by the generic software value, which is used when the proposal is intended to work with any software, rather than targeting and exploiting the peculiarities of a specific application domain.
How? The taxonomy in Table 3 enables classifying eliciting the publication trends for the selected studies. More precisely, it enables indicating the publication year for each selected study, as well as its publication type (i.e., book chapter, conference paper, or journal article).
2.5. Replication Package
To enable repeating our study, as well as verifying our findings, we publicly released a replication package on Zenodo (Danushi et al., 2025). The replication package includes the intermediate artifacts produced during the search and selection of studies to be considered, as well as their final classification. The selected studies can instead be accessed by relying on the bibliographic information provided by the references available in Table 2.
3. Overviewing the State-of-the-Art Solutions
This section overviews the whole corpus of our SLR, which constitutes the state of the art in the field of designing and developing environmentally sustainable software. Particularly, we hereinafter present the surveyed articles by dividing them as per the What? classification from Section 2.4, namely sustainability guidelines (Section 3.1), reference models (Section 3.2), measurement solutions (Section 3.3), and software improvement techniques (Section 3.4). To support the reader, the classification of the selected studies is graphically recapped in Table 4.
| Who? | What? | Why? | Where? | When? | How? | ||||||||||||||||||||
| indirectly | directly | ||||||||||||||||||||||||
| Ref. |
end user |
IT operator |
product owner |
sust. engineer |
software architect |
software developer |
software provider |
proposal |
white-/black-box |
TRL |
measure |
reduce |
minimise |
measure |
reduce |
minimise |
target app. domain |
requirement analysis |
design |
development |
integration & testing |
roll-out |
publication type |
publication year |
[…]
3.1. Sustainability Guidelines
In this section, we focus on the portion of the surveyed corpus that proposes sustainability-related guidelines to design and develop sustainable software. We start by illustrating those studies that focus on indirectly reducing the carbon footprint of running software. We then move to those directly considering the carbon footprint of software, and finally discuss those works that follow a hybrid approach. We organise our discussion by target application domain.
Indirect. For generic software, Kharchenko and Gorbenko (Kharchenko and Gorbenko, 2013) provide a set of good practices in the form of procedural steps intended for sustainability engineers to implement green management of system components. Such guidelines account for different aspects of system components (e.g., threats, fault-tolerance) and enable identifying existing pathological and evolutionary chains among those aspects. From the analysis of such chains it is possible to determine how to minimise the system’s resource (and energy) consumption, while improving a set of customisable quality metrics (e.g., reliability, availability, security, resilience).
On a different line, addressing generic software architects, Ponsard et al. (Ponsard et al., 2018) conduct an experience report of an industrial case study about building and putting in operation a shared pick-up service. They show how to conduct the requirement analysis phase by also considering sustainability goals and their intertwining with the functional requirements set by other stakeholders of a software system. By distinguishing between green by software and green in software requirements, they take implementation and architectural decisions at different levels (e.g., choice of efficient programming language and algorithms, use of lightweight virtualisation) and define key performance indicators to monitor for improving future releases.
Moving to approaches that consider HPC as their application domain, Radersma (Radersma, 2022) lists a set of good practices for software developers to implement more environmentally sustainable and maintainable software artefacts, e.g., by adopting green programming languages, avoiding unnecessary data accesses in favour or computing results on-the-fly, employing state-of-the-art HPC hardware, and relying on cluster power management software to optimise energy consumption. Considering instead an experience report on ML, Brownlee et al. (Brownlee et al., 2021) rely on exhaustive and genetic search to (sub-)optimally tune the hyperparameters of multi-layer perceptron networks. From a software developer’s perspective, they investigate the relations among different configurations, accuracy and energy consumption in both the training and inference phases on a set of benchmark networks. As a result, in their benchmark, they reduce energy consumption by 30–50% with a decrease in accuracy below 1%.
Direct. Still in the ML domain, Pan and Rajan (Pan and Rajan, 2022) are the sole to discuss good practices for software architects to directly reduce carbon emissions produced by convolutional neural networks. Instead of training convolutional neural networks from scratch, they recommend exploiting model decomposition – having each module capable of classifying a single output class – together with transfer, one-shot, and few-shot learning to favour reusability and replaceability and (consequently) reduce the carbon emissions associated with model (re)training.
Hybrid. In collaboration with an industry partner and targeting software developers, Fernandez and Lago (Condori-Fernández and Lago, 2018) identify a set of domain generic green strategies and link them to the prioritisation of software quality requirements (e.g. usability, maintainability). Their aim is threefold: to raise people’s awareness of software sustainability, to make service operations environmentally sustainable, and to support sustainable development processes, by emphasising the need to consider sustainability throughout the software design phase and beyond. Similarly, Shenoy and Eeratta (Shenoy and Eeratta, 2011) suggest that quality and sustainability should be considered equally important throughout the whole software lifecycle. They propose a set of improvement practices for software developers to drive green software development.
Still targeting generic software, Mehra et al. (Mehra et al., 2023) report on their experience of using a static analysis tool to identify code inefficiencies and act upon them suitably. Their experience report, clearly intended for application developers, showcases the potential of static analysis in detecting energy-inefficient patterns, assessing their impact, and providing remediation strategies. They indeed successfully refactor a large application codebase from an industrial use case. Siegmund et al. (Siegmund et al., 2022) analyse the relationship between software configuration and energy consumption. They propose ML-based code analysis methods to identify energy and carbon inefficiencies in software, from the developer’s perspective. On the other hand, they enable software users to select energy-optimised configurations for their software according to a black-box model. Their model accounts for non-monotonic energy consumption in the configuration space, which enables relying non-linear functions to model the system’s behaviour.
Building upon the work by Shenoy and Eeratta (Shenoy and Eeratta, 2011), Chauhan and Saxena (Chauhan and Saxena, 2013) propose the Green Cloud Framework (GCF) as a set of best practices to handle software development lifecycle in cloud computing settings. Their work accounts for recommendations from organisations like Greenpeace to use clean and renewable energy sources in software development to make software development energy-aware and carbon-aware. Specifically, GCF is intended for software providers, developers and end-users, to make Service Level Agreements (SLAs) include guarantees of sustainability metrics (i.e. carbon emissions and energy consumption). Differently, (Shenoy and Eeratta, 2011) mainly focuses on software developers, without envisioning the key role of SLAs between application providers and end-users.
Ahmed (Ahmed, 2020) introduces a white-box technique for software developers to study cloud software via direct access to the source code. He proposes a new step to optimise the software development lifecycle, named parameterized development phase, between the design and implementation phases, useful to define parameters related to the used programming language and hardware architecture, and their energy consumption. This new step is useful to identify margins for improvement as early as possible in the design phase to allow for code tuning in a cost-effective manner.
Speaking of ML, Patterson et al. (Patterson et al., 2022) discuss the energy consumption and carbon footprint of machine learning training and inference, also considering embodied carbon emissions. They eventually propose the so-called 4M best practices for software developers, which involve choosing greener energy locations for ML workloads (Map), using the latest possible hardware (Machine), relying on cloud resources to reduce the environmental footprint (Mechanisaton), and exploit state of the art lightweight neural networks (Models). Similarly, Fu et al. (Fu et al., 2021) provide good practices for software developers to compute carbon emissions costs for the search phase of a computer vision architecture based on deep learning over its entire lifecycle.
3.2. Reference models
In this section, we focus on the largest part of our surveyed corpus, which proposes reference models for the design and development of carbon-aware applications. As in the previous section, we first report on contributions that indirectly reduce carbon emissions, and then move to those that focus on direct carbon reduction. Finally, we discuss hybrid proposals.
Indirect. Sharma et al. (Sharma et al., 2015) and Mohankumar and Kumar (Mohankumar and Kumar, 2016) introduce a reference model for software developers, by relying and extending the software lifecycle model of (Kneuper, 2018). Sharma et al. (Sharma et al., 2015) consider an iterative software development lifecycle made of six stages, i.e., requirement analysis, green analysis, design, implementation, testing, and maintenance. The green analysis phase cyclically connects to all others and acts as a checkpoint entity to monitor and continuously enhance the sustainability of all other phases according to set guidelines. Also following an iterative approach, Mohankumar and Kumar (Mohankumar and Kumar, 2016) only consider a four-stage lifecycle consisting of requirements analysis, design, coding and testing. For the requirements analysis, they propose to use questions, guidelines, and generic templates to match customers’ needs with sustainable technology choices, use well-timed metrics data for the purpose, and formal methods to avoid ambiguities. Based on requirements specification, developers should write the documentation concerning green and sustainable templates for designing all application components. Then, throughout the implementation phase they should always adopt best practices, e.g. avoid unnecessary classes, closing database connections after use, avoid stand-alone installation in favour of on-demand remote access.
The reference model by Kern et al. (Kern et al., 2013) goes towards defining an eco-labelling to drive software developers in the design and development phases. It can be applied by analysing which tasks or functions of the software are most frequently used and then designing case studies accordingly, also considering servers’ energy consumption. This way of proceeding allows developers and end-users to compare and contrast different software implementations by determining the most sustainable one. Subsequent proposals by Kern et al. are also aimed at IT operators (Kern et al., 2013) and for sustainability engineers and project owners (Kern et al., 2018).
Direct. To reduce carbon footprint and to raise green and sustainable awareness in generic software, Lami et al. (Lami et al., 2014, 2013) define a methodological approach for product owners. They use such an approach to derive sustainability indicators for the whole lifecycle of a software project and use them to populate bi-dimensional matrices, known as hierarchical software templates. Such templates are used to explore three levels of environmental effects, viz. direct, indirect, and systemic, also associating direct effects with their measurements and success factors. Last, in compliance with ISO/IEC 33000, Lami et al. (Lami et al., 2014) define the assessment of six process sustainability levels, viz. incomplete, performed, managed, defined, measured, and optimised.
Considering instead the peculiarities of cloud applications, Atkinson et al. (Atkinson et al., 2014) propose a reference model that is intended for several stakeholders such as product owners, software providers, developers, and end-users. The model focuses on the specification of SLAs to integrate energy costs and environmental impact (e.g., usage of renewable energy sources) among the considered QoS parameters. They propose adopting different UML modelling to better communicate with all stakeholders by relying on structural, functional, and behavioural views. With an eye to economical sustainability, their proposal also models different pricing policies of the software offer.
Hybrid. All studies on hybrid reference models are intended for the generic software domain with a black-box approach. For example, Tee et al. (Tee et al., 2014a, b) define a reference model for both software developers and IT operators, by proposing a procedural method for making software development lifecycle more sustainable. They aim at gathering, assembling and storing of the information for the knowledge management of a green and sustainable software development lifecycle. They collect information on existing findings, processes and activities, and techniques in software development, and propose a knowledge management model and step-by-step guidelines to make software sustainable. They then map techniques to actionable tasks to minimise power consumption and carbon footprint while developing software. Following up on the above, Tee et al. (Tee et al., 2014a) extend the knowledge-management model into a green-management model to be shared among all stakeholders to build an action-oriented and collaborative community of practice.
Jayanthi et al. (Jayanthi et al., 2021) and Dick et al. (Dick et al., 2013) provide a reference model for software developers, sustainability engineers, and end-users by defining green and sustainable software development lifecycle steps integrated within Agile operations. In particular, Jayanthi et al. (Jayanthi et al., 2021) argue that organisations should establish a sustainability department that includes quality assurance engineers for the requirements analysis phase, user experience designers for the design phase, certified green computing developers for the coding and testing phase, and managers for the overall success of the sustainable software deployment. Conversely, Dick et al. (Dick et al., 2013) add a specific sustainable development stage to the software lifecycle model. They associate SCRUM activities to their driving role among sustainability engineers and software developers, and embed sustainability reporting and retrospective in the Agile development model.
On a similar line, extending their previous effort (Lami et al., 2014), Lami and Buglione (Lami and Buglione, 2012) provide a reference model for project owners to enact measurements by considering a cause-effect diagram analysis on the infrastructure, people, software process and product. The model has to create a suitable measurement plan that integrates all needed information elements to specifically address the sustainability improvement throughout software development. They add sustainability to project management as a first-class citizen like quality, time, and budget. With an akin objective, Ibrahim et al. (Ibrahim et al., 2021, 2022) provide a reference model for both software developers and end-users. They focus holistically on software product development, as well as the related hardware disposal and recycling. Their approach also relies on protocols for (semi-)structured interviews with involved stakeholders to collect data useful for process mining from a sustainability viewpoint. The proposed model aims at achieving resource and energy efficiency to pursue long-term improvement of software sustainability, with a particular focus on reducing waste (e.g., over-requirements, miscommunication, rework, and lack of knowledge). Ibrahim et al. (Ibrahim et al., 2021) also showcase the proposed approach over real industrial use cases.
Raisian et al. (Raisian et al., 2021) propose green and sustainable software measurements as a reference model for software developers that aim at targeting sustainability by considering the environmental dimension (i.e., energy and carbon footprint in processing and storage), the economic dimension (i.e., usability, available time and costs), and the social dimension (i.e., support tools, human well-being, performance). To achieve the above, they aim at identifying suitable hierarchical measurements that are to be evaluated, documented, and assessed for continuous software enhancement. Still with a hierarchical approach to continuous software improvement, Penzenstadler and Femmer (Penzenstadler and Femmer, 2013) provide a reference model definition based on different levels of granularity, for both software architects and sustainability engineers. After defining sustainability metrics and their targets, the proposed model foresees matching software lifecycle activities to their (positive/negative) impacts on energy or carbon emissions, and identifying responsible personnel for each activity.
Becker et al. (Becker et al., 2016) propose a reference model that allows product owners, software providers, developers and sustainability engineers to compare two equivalent software systems designed and implemented with and without considering sustainability, respectively. The model aligns the sustainability goals with those of the system, defines the project scope and system boundaries for any stakeholder, makes the risk analysis for each sustainability dimension, and finally, implements requirements and specifications based on templates including checklists for sustainability criteria and standards compliance in all the interrelated dimensions. Lago and Jansen (Lago and Jansen, 2010) propose a reference model that supports the iterative and incremental way of operating for software developers and end-users. The model is used to drive the implementation of two Web services, i.e., the Environmental Strategies as-a-Service and the Green Metrics as-a-Service, useful to make the environmental impact of software services visible by continuously monitoring it, giving feedback and tailoring software based on the available data to reduce energy and carbon footprints.
We conclude by presenting the GREENSOFT reference model, originally defined by Naumann et al. (Naumann et al., 2011), which led to many other research efforts. Following a holistic approach, GREENSOFT is intended for all stakeholders, i.e.,software architects, developers, IT operators, and end-users. The model seconds the Agile software development model and considers first (green-IT), second (green-by-IT) and third order (systemic) effects related to designing, implementing, using and decommissioning software products. It foresees the definition of ad hoc sustainability criteria and metrics, the analysis of process models for the development, administration, purchase and use of software, and the adoption of recommendations and tools for all stakeholders. Agile sprints are enhanced by including sustainability reviews and retrospectives that focus on measured metrics and criteria, monitored via a sustainability journal throughout requirements elicitation, design, implementation and testing. These steps are mentioned by Dick et al. (Dick et al., 2013) when considering the sustainable Agile development stage.
Retaking the process assessment model of GREENSOFT, Johann et al. (Johann et al., 2011) propose a reference model to develop sustainable software products, while mitigating negative impacts and amplifying positive impacts based on a knowledge base shared among different stakeholders. Their proposal includes a first implementation of recommendations and tools in the form of checklists and suggestions. Dick et al. (Dick et al., 2010) also extend GREENSOFT by relying on the definition of checklists and best practices realised through stakeholder interviews, with a focus on software development and project management. Particularly, they highlight how sustainability decisions taken at a set step of the software lifecycle only show their impact during the next one.
Later, Mahmoud and Ahmad (Mahmoud and Ahmad, 2013) re-consider the GREENSOFT model and discuss how software itself can be used as a tool to aid and promote fine-grained green computing for operating systems, frameworks, performance monitoring, programs written for energy management purposes, and virtualisation. Their model also accounts for fault-tolerance and failure management. Finally, following GREENSOFT and the ISO/IEC 12207 standard,444Systems and software engineering — Software lifecycle processes, International Organization for Standardization, 2008 Naumann et al. (Naumann et al., 2015) define criteria to identify software quality aspects as per a process-centric model, inherent both to the final product and the implemented process, through lifecycle assessment. Their model addresses software developers, sustainability engineers and end-users and is given in a full-fledged Agile version following the whole lifecycle as well as in a lightweight version focussing on process assessment and sustainability reviews and previews.
3.3. Measurement Solutions
This section focuses on the portion of the analysed corpus about measurement solutions for software sustainability, which take the form of estimate techniques and assistive tooling. Their aim is to measure or estimate carbon emissions related to software design and development so as to make informed choice throughout the software lifecycle.
Indirect. Starting from those approaches that indirectly consider carbon emissions, Noureddine et al. (Noureddine et al., 2012) develop a real-time software energy monitoring framework intended for both software developers and end-users. Such solution reports estimates on energy consumption by relying on a service-oriented architecture consisting of a monitoring sensor, a database, and a lifecycle monitoring module. Specifically, the individual components are isolated and managed using start/stop options, thus providing flexibility while continuously monitoring the deployment hardware (i.e., processors, network interfaces). The approach is illustrated over different programs written in different programming languages.
Focussing instead on the HPC domain, Kipp et al. (Kipp et al., 2011) introduce an estimate technique intended for sustainability engineers. They propose Green Performance Indicators as a supplement to Key Performance Indicators encompassing factors such as the use of virtual/real IT resources, the service lifecycle, approximate environmental impact, and organisational considerations. They consequently assess sustainability at different levels of a system architecture - from the abstract view to the business logic and up to the entire IT service centre. The methodology involves monitoring clusters of metrics for energy-intensive applications by first analysing single jobs and observing the relation between energy consumption and execution time for each run. Such data is then input to a constraint-based formulation to optimise resource usage and energy consumption.
Direct. Sharma et al. (Mehra et al., 2022) devise an assistive tooling for software developers to assess the carbon efficiency of their projects throughout the development lifecycle. This tool provides a standard set of metrics to evaluate the energy consumption impact of different design decisions at each stage of the project. The approach is based on a series of questions related to project characteristics, e.g., programming languages, used third-party libraries, cloud providers, technologies, processes, and teams’ organisation. Ultimately, the tool provides an overall sustainability score and recommendations to further improve software sustainability. To also promote responsibility, awareness and transparency, Kern et al. (Kern et al., 2015) develop a model- and tool-based calculation method intended for software developers and product owners to measure the carbon footprint of their products according to international standards. The strategy involves collecting data throughout the software development lifecycle to continuously monitor carbon emissions.
Hybrid. This last section contains a large number of surveyed studies, scattered across a plethora of different domains. Starting from generic software, Pa et al. (Pa et al., 2017) propose an assistive tool to help developers in understanding and implementing sustainable software by estimating its carbon footprint, also based on its energy consumption. The prototype features online accessible web-based dashboards to visualise sustainability reports and suggestions for defining best trade-off solutions between database organisation, network usage and human resources. The tool can also track historical data to monitor project evolution over time. Simon et al. (Simon et al., 2023) provide a comprehensive estimate technique based on graph modelling regarding the environmental impact of software lifecycle for project owners. They have nodes representing lifecycle stages and edges describing different environmental impacts based on direct observations. Lifecycle assessment follows a lightweight bottom-up approach to identify stakeholders’ responsibility and actionable insights and to identify hot spots for potential improvement. Still targeting project owners, Moshnyaga (Moshnyaga, 2013a, b) analyses the total energy consumption associated with the software throughout its entire development lifecycle and evaluates its contribution to carbon emissions, also accounting for hardware-specific emissions. The proposed methodology also relies on rapid prototyping to build software following an iterative three-step process consisting of business model definition, risk analysis (revision), and development plan (update). The same authors also tailored their proposal specifically targeting mobile applications (Moshnyaga, 2013c), along with both project owners and software developers.
Speaking of the ML domain, based on Mehra et al. (Mehra et al., 2022), Sikand et al. (Sikand et al., 2023) present a standard set of metrics for measuring the energy efficiency of artificial intelligence systems throughout the development lifecycle. They also call attention to harmful aspects associated with red AI, e.g., the exponential growth in data, model complexity, large models, increased number of experiments, and the relentless pursuit of higher performance. Extending their previous effort (Kipp et al., 2011), Kipp et al. (Kipp et al., 2012) instead focus on HPC by introducing Green Performance Indicators to complement the traditional KPIs, by enabling a comprehensive evaluation of the environmental impact and sustainability classification of IT service centres by examining various hierarchical system levels, from the application level to the middle-ware and up to the infrastructure.
3.4. Software Improvement Techniques
We last explore the portion of our surveyed corpus that proposes specific improvement techniques to design and develop sustainable software, e.g.,decision-making, model-driven, static analysis, testing, optimisation, search-based software engineering, and distributed protocols.
Indirect. Kocak and Alptekin (Koçak and Alptekin, 2019) propose a multi-criteria decision-making technique for generic software, as part of the analysis and design of carbon efficient software. Their technique aims at balancing between the needs of end-users, who want energy-efficient software at a low cost, and software providers, who offer products at a maximum quality level. The proposal addresses the conflicting interests by determining demand and supply functions for decision-making entities that are considered rational, intelligent, and having a comprehensive plan of action for any situation that may arise. The proposal also discusses the implementation of a strategic interaction based on the interdependencies of software product properties and attributes. They also align their proposal with the ISO/IEC 25010 standards, including attributes such as functional suitability, reliability, performance efficiency, and usability. The authors then propose a method to translate these requirements into terms of price, quality, and utility.
To ensure software functionality compliance and measure energy consumption, Wedyan et al. (Wedyan et al., 2023) propose a white-box analysis technique for testing purposes, to software developers. They advocate the use of static analysis to construct call graphs for each class, with methods denoted by nodes, and control between methods denoted by edges. They then associate an estimate of energy consumption per computation unit and recommend considering cyclomatic complexity, which is a measure of software complexity based on the number of decision points and lines of code involved. Their testing framework ultimately reports both the energy consumed and the success of the test itself. Similarly, Mancebo et al. (Mancebo et al., 2021) propose a white-box analysis technique for testing purposes to software developers, IT operators and end-users. Their prototype framework is capable of gathering energy consumption data during software execution, as well as of processing them to generate statistics and reports for comparison and analysis. They provide users with a conceptual framework for green and sustainable software measurement, a methodological guide to conduct energy consumption measurements, and a prototype tester.
Focussing on mobile applications, Le Goaër (Goaër, 2021) provides a code scanning tool for the Android Studio IDE. The tool performs white-box static analyses to identify green code smells and associated potential energy optimisation. The scanning tool is designed for software developers who use the Android Lint framework. Although individual energy savings may be small for single Android devices, the cumulative impact across billions of devices is significant. Different green bugs are matched to their severity, priority, and category, along with candidate single-click fixes (e.g., rely on dark UI, data compression, timeout wave locks). Still in the mobile landscape, Beghoura et al. (Beghoura et al., 2014) instead propose a black box analysis for software developers that relies on a supervised vector machine model to create specific energy consumption models for different types of devices. They investigate the energy consumption patterns of hardware when running specific applications, accounting for various factors such as hardware configuration, product design, and battery ageing. More specifically, the study aims at understanding non-linear energy consumption components, such as changes in CPU power consumption as a function of frequency and voltage.
Moving to the field of blockchains, Lasla et al. (Lasla et al., 2022) introduce a distributed protocol, that defines a green proof-of-work algorithm. The implementation comprises two rounds. The first involves all network nodes in the mining process, while the second involves a smaller set of nodes, called runners-up. The energy consumption is reduced since the same energy spent to mine one block is also used to select a small number of miners to exclusively mine the next block. The protocol also considers and balances among usual blockchain tradeoffs such as fork frequency, security and decentralisation level.
Direct. Only the proposal by Karunakaran and Rao (Karunakaran and Rao, 2013), intended for project owners, provides a model-driven technique that solely considers the carbon footprint of software. They propose to exploit coloured Petri Nets that embed time constraints and hierarchical implementation to devise and configure optimal policies of a waterfall software development lifecycle. Their model accounts for machine, software and resource levels to simulate programming, testing and debugging of software projects and estimate the associated carbon emissions.
Hybrid. Hoesch-Klohe et al. (Hoesch-Klohe et al., 2010) provide a model-driven technique which leverages business process models for their capability to capture the relations between resources and activities, specifically focusing on carbon emission impact, through textual assertions. Their technique involves identifying all process designs, ordering them based on their abstraction level, and then systematically determining a suitable association to computing resources accounting both for carbon emissions and energy consumption. Lago (Lago, 2019) devises a decision-making technique intended for software architects, based on architectural designs and assessments for sustainability. In the specific, decision maps capture the impacts of sustainability dimensions as immediate, enabling, and systemic, and categorise their effects as positive, negative, or undecided. Decision maps are used to explore the software design space by addressing multiple sustainability perspectives at the same time; they can be used to drive the software design decision process, while refining it iteratively.
Focussing on cyber-physical systems, Heithoff et al. (Heithoff et al., 2023) instead present a model-based approach for project owners and software developers. By relying on digital twins and digital shadow systems, they monitor the sustainability of the actual real-world software. Digital twins are designed to be smaller in size than the original system and allow for faster analyses through simulated behaviour. Their prototype also enables monitoring energy consumption peaks in comparison to current workloads and predefined levels. With cloud computing developers as their target, Jebraeil et al. (Jebraeil et al., 2017) propose a model-driven technique called g-UML. They extend UML diagrams with contextual information about data centre and network operations, i.e., power generation, cooling and servers consumption, data communication footprint. This enables the efficient identification of energy-related issues and the organisation of power-related components into an architectural view. Resulting gUML diagrams can assist software developers in writing energy-efficient software, enabling collaboration between hardware and software architects. Addressing software developers, Fakhar et al. (Fakhar et al., 2012) combine a white-box static analysis technique for energy-efficient software compilation with a black-box analysis technique for effective task scheduling. The compiler translates source code in an energy-efficient format, optimises binary code through green and sustainable strategies, and provides hints for energy-efficient coding. On the other hand, the scheduler utilizes a greedy algorithm to efficiently make use of cloud resources, selecting and allocating machines based on user requirements while aiming at using as many processing cores in a machine as possible.
Moving from cloud to edge computing settings, Zhang et al. (Zhang et al., 2023) present an optimisation technique tailored for software providers operating in a network that offers smart services. They propose the use of an evaluation framework that combines dynamic energy trading with task allocation optimisation for federated edge intelligence networks to reduce carbon emissions throughout the software lifecycle. At design time, they mainly focus on computing carbon emissions related to data-generating devices. At development time, they account for emissions due to pre-training, testing, parameter selection and optimisation, based on specified service requirements and usage scenarios. At deployment time, they address model distribution, while, at disposal time, they focus on generalisation, modification, and transfer issues.
In line with Le Goaër (Goaër, 2021), Le Goaër et al. (Goaër, 2023; Goaër and Hertout, 2022) present a white-box static, computer-aided, customisable code inspection tool for Android software developers, named ecoCode. This tool is designed to identify and refactor code structures that are inefficient in terms of energy consumption (thus impacting on carbon emissions). By emphasising energy efficiency as a quality attribute, ecoCode aims at promoting the concept of green and sustainable code, akin to the way security and maintainability are enforced in clean code. Particularly, ecoCode performs a scan of project repositories to identify environmental issues, which are then displayed on a web dashboard.
To reduce re-training costs in ML, Wei et al. (Wei et al., 2023) propose a machine learning technique for software developers. Specifically, they suggest using model compression through a quantization method for the tasks of code generation (from natural language to code), summarisation (the counterpart, from code to natural language), and type inference. This technique involves representing model parameters with fewer bits, making the model smaller in size so as to increase the speed of operations. Empirical results validate the effectiveness of this approach on multiple code-specific tasks, demonstrating that it is green and sustainable while preserving accuracy and robustness compared to the full precision counterpart model.
Finally, in the domain of blockchains, Alofi et al. (Alofi et al., 2023, 2022) propose a search-based software engineering technique for addressing energy and carbon consumption concerns. The proposal is intended for software architects and sustainability engineers and employs evolutionary meta-heuristics to identify subsets of miners that maximise trust levels in the system, maintain decentralisation and reduce energy consumption and carbon emissions. They propose to implement the system as an autonomic MAPE-K loop that ensures enactment of taken decisions.
4. Answering our 5W1H Questions
This section aims at answering to our 5W1H questions on the design and development of environmentally sustainable software (Section 2.1). More precisely, by zooming out and considering the overall classification given by Table 4, we aim at identifying trends and coverage on who are the targeted stakeholders (Section 4.1), what methods are proposed to design and develop environmentally sustainable software (Section 4.2), why (Section 4.3), where and when the proposed solutions are applied (Sections 4.4 and 4.5, respectively), and how the selected studies have been presented to the scientific community (Section 4.6).
4.1. Who?
The who of the design and development of environmentally sustainable software aims at identifying who are the stakeholders targeted by the state of the art solutions available for the topic. Figure 2 lists the stakeholders considered by the selected studies, as well as their frequencies.555In this and the following subsections, we refer to the frequency of a given classification in the selected studies, measured by the number of studies classified as such.
From Figure 2, there is an evident unbalance between design- and development-related stakeholders. Indeed, software developers are by far the most frequent stakeholders, who are targeted by the vast majority of selected studies (i.e., 45 out of the 65 selected studies). The frequency of software developers is also significantly higher than that of software design-related stakeholders (i.e., software architects and sustainability engineers), even when added together. Given that our focus is on the design and development of sustainable software, we would have expected a more balanced distribution among design- and development-related stakeholders. In this perspective, and to avoid putting all the burden of environmental sustainability on software developers, there is a need for more solutions targeting software architects and sustainability engineers, who should be better supported in the design of environmentally sustainable software.
Other interesting observations come from the coverage of stakeholders not directly involved in software design and development, like end users, product owners, IT operators, and software providers. While the frequency of IT operators may be justified by the emergent wave of DevOps practices (Azad and Hyrynsalmi, 2023), the frequencies of the other stakeholders seem to suggest that the design and development of environmentally sustainable software naturally involves more stakeholders than just those involved in the design and development of software. This holds especially for end users and product owners, whose frequencies are almost the same as those of design-related stakeholders, since they can concretely impact on the design and development of environmentally sustainable software. Indeed, environmental sustainability can be a discriminant for the software chosen by end users (Jayanthi et al., 2021; Lago and Jansen, 2010; Atkinson et al., 2014; Mancebo et al., 2021). Additionally, the choices taken by product owners along the software process can affect the environmental sustainability of the software itself (Lami et al., 2014; Moshnyaga, 2013a; Sikand et al., 2023; Moshnyaga, 2013c). The above are just first insights on how different stakeholders impact on the design and development of environmentally sustainable software, even if not directly participating in software design and development themselves. Such insights shed light on the need for further investigation on the roles, contributions, and support needed by all the possible stakeholders for the design and development of environmentally sustainable software.
| Highlights | |
| Software developers are by far the most frequent stakeholders for the existing solutions to support the design and development of environmentally sustainable software. | |
| More support should be provided to software architects and sustainability engineers for the design of environmentally sustainable software. | |
| The design and development of environmentally sustainable software involves multiple different stakeholders, whose roles, contributions, and needed support should be further investigated. | |
4.2. What?
We hereby consider what are the state of the art solutions available for designing and developing environmentally sustainable software. More precisely, we aim at providing an overall view of the available proposals, whether they directly inspect source code, and their overall TRL, based on the classification of selected studies provided in Table 4.
Figure 3 displays the frequency of proposals in the selected studies, clearly outlining a predominance of reference models, proposed in 23 out of the 65 selected studies. Of these, six studies focus on GREENSOFT, a reference model first proposed by Naumann et al. (Naumann et al., 2011) and then rivisited and extended in five subsequent studies (Dick et al., 2013; Johann et al., 2011; Dick et al., 2010; Mahmoud and Ahmad, 2013; Naumann et al., 2015). The remaining efforts propose different reference models, however still sharing some commonalities with GREENSOFT and with each other. Whilst the number of reference models witnesses the need for clarifying what is environmental sustainability for software design and development, the lack of a commonly accepted reference model calls for standardisation efforts along this line.
The existing proposals for designing and developing environmentally sustainable software then mainly focus on providing a first support to the purpose, namely sharing guidelines in the form of good practices or providing estimate techniques to assess the environmental sustainability of the software under design/development. The rest is then scattered among different attempts (especially, software improvement techniques) to support the design and development of environmentally sustainable software, however without any concrete proposal emerging as the most promising or used for the task. Further research is, therefore, needed to establish which techniques can effectively support designing and developing environmentally sustainable software.
On another front, Figure 4a shows that the vast majority of the existing proposals adopt a black box approach. Namely, most of the existing proposals do not need the source code of software to be available. This is mainly motivated by the above distribution of proposals, since reference models and good practices (which are the most frequent categories) typically abstract from the source code of a software. Another motivation comes from the black-box software improvement techniques (Karunakaran and Rao, 2013; Heithoff et al., 2023; Beghoura et al., 2014; Jebraeil et al., 2017; Alofi et al., 2022, 2023; Tee et al., 2014a; Wei et al., 2023; Zhang et al., 2023; Lago, 2019; Lasla et al., 2022; Koçak and Alptekin, 2019), which choose a black-box approach to increase their applicability, namely to work also with software whose sources are not available, e.g., for privacy or legacy reasons. At the same time, the existing white-box proposals show concrete potentials to enhance the environmental sustainability of the software under design/development, as witnessed, e.g., by the green code smells detected by the ecoCode project (Goaër, 2021, 2023; Goaër and Hertout, 2022). Following this line, it may be worthy to consider further investigating the potentials of white-box analysis, as well as finding trade-offs between proposals’ applicability and enhancements in terms of sustainability.
Finally, Figure 4b displays the overall low TRL of the state of the art on designing and developing environmentally sustainable software. Most of the studies feature a TRL of 2 or 3, which – together with TRL 1 – cover around 66% of the selected studies. Among those, only 21.5% of the selected studies went beyond observing the basic principles and formulating the technology concepts, by proposing an experimental proof-of-concept. 32.3% of the selected studies went a step higher, by validating their proposal in laboratory and, therefore, reaching TRL 4. Only the experience report by Mehra et al. (Mehra et al., 2023) reached TRL 5, being it a report on an industry-relevant case study. Overall, the above described distribution of TRLs witnesses a need for increasing the TRL of what is available to support the design and development of environmentally sustainable software, especially if we wish to support the software industry in this task.
| Highlights | |
| Most of the proposals are reference models, witnessing the need for clarifying what is the environmentally sustainable software design and development. | |
| Existing reference models differ one another, calling for standardisation efforts. | |
| Black-box analyses are the most used, motivated by their applicability. White-box analyses show potentials and should be investigated further, along with trade-offs between applicability and sustainability enhancements. | |
| The overall TRL is low and should be increased to favour the adoption of existing solutions by the software industry. | |
4.3. Why?
The why question aims at identifying the sustainability goals in the selected studies, namely whether they support the design and development of carbon-efficient software by directly or indirectly considering its carbon footprint. As shown in Figure 5, the most frequent trends are reducing the carbon footprint of software through a hybrid approach, namely directly considering a software’s carbon footprint, as well as considering the indirect effect of its energy consumption on such footprint.
This is motivated by the types of proposals targeting such goals, i.e., reference models and sustainability guidelines provide high level information, which can effectively be exploited to reduce the energy consumption or carbon footprint of software – but not enough detailed to measure or minimise them.
There are also solutions proposed to enable directly/indirectly measuring carbon emissions, which are among the goals of ten and seven of the selected studies, respectively. The underlying idea is in line with that of the studies targeting the reduction of carbon emissions. More precisely, the solutions aimed at directly/indirectly measuring carbon emissions are intended to enable comparing the carbon emissions before/after updating a software, with the ultimate goal of assessing whether any applied or planned change contributes to reducing them (Patterson et al., 2022; Kern et al., 2013; Lasla et al., 2022; Kipp et al., 2011).
Only three selected studies (Alofi et al., 2023, 2022; Zhang et al., 2023) pursue the goal of minimising the carbon footprint of software, doing this with a hybrid approach. Such a low number is not only motivated by the type of proposals, as highlighted above, but also by the inherent complexity of optimising by considering the multiple different aspects that influence the carbon footprint of the software under design/development (Mancebo et al., 2021; Siegmund et al., 2022; Wedyan et al., 2023; Tee et al., 2014a; Mehra et al., 2022). Indeed, Alofi et al. (Alofi et al., 2022, 2023) and Zhang et al. (Zhang et al., 2023) pursued their optimisation goals by narrowing the search space by relying on the context given by blockchains and ML, respectively. These are quite recent efforts, which anyhow showed the feasibility of minimising a software’s carbon footprint. Further investigation is needed along this line, possibly also considering, e.g., heuristics to achieve sub-optimal minimisation, if the holistic minimisation would get too complex to achieve.
Other interesting observations come from Figure 6, which shows that only seven out of the 65 selected studies consider only the carbon footprint of software. Indeed, most of the selected studies pursue their carbon efficiency goals with an indirect or hybrid approach (in 18 and 40 of the selected studies, respectively). In both cases, the carbon footprint is estimated starting from energy consumption while pursuing the goal of measuring/reducing/minimising a software’s carbon footprint. This is mainly motivated by an assumption shared by the selected studies considering energy consumption, namely that reducing the energy consumption of software also reduces its carbon footprint. However, as demonstrated by Aiello et al. (Aiello et al., 2024), energy mixes play a crucial role to relate energy consumption with carbon footprint, as different energy mixes have different associated carbon intensities. This may negatively impact on the effectiveness of considering energy consumption to indirectly reduce carbon emission. For example, when a software is powered with clean and renewable energy, there may be no need for reducing its energy consumption, as this may not significantly change its carbon footprint (Forti et al., 2024). This calls for further research along the line of directly considering energy mixes and carbon footprint, rather than as an indirect consequence of energy consumption.
| Highlights | |
| The most pursued goals are reducing carbon emissions. Their minimisation can anyhow be achieved by relying on the application domain to narrow the search space of the minimisation problem. | |
| Reducing energy consumption to indirectly reduce a software’s carbon footprint may not be enough, as different energy mixes may result in different carbon emissions. This calls for solutions directly considering energy mixes and carbon footprint as well. | |
4.4. Where?
We hereby analyse the application domains, if any, targeted by the existing solutions for the design and development of environmentally sustainable software.
Figure 7 displays the frequency of application domains in the selected studies, outlining that the vast majority of existing solutions target generic software – rather than application domain-specific software.
This is mainly motivated by the fact that sustainability guidelines, reference models, and measurement solutions are given as high-level, broadly-applicable approaches to reduce the carbon emissions of software, as it can be observed from the heatmaps in Figure 8. The applicability and extension of such generic approaches to specific application domains is, therefore, deserving further investigation, as they might require some adaptation for getting applied to application domain-specific software. The same does not hold for the available software improvement techniques, as most of them target specific application domains (Figure 8), rather exploiting the peculiarities of application domain-specific software for reducing its carbon emissions.
Going beyond generic software, the most targeted application domains are cloud computing, ML, and mobile applications, followed by blockchain and HPC. Only single spotted contributions, instead, target the domains of computer vision, cyber-physical systems, and edge computing. Most of the efforts are indeed geared towards reducing the environmental impact of computationally intensive application domains, which often come with significant energy consumption and carbon footprint (Pan and Rajan, 2022; Zhang et al., 2023; Kipp et al., 2011). This holds for all application domains mentioned above, but for the case of mobile applications, where the focus is more on limiting the energy draining of mobile devices. The idea is essentially that, even if a small amount of energy is saved on a mobile device, such an energy save have to be summed over billions of devices, and this can contribute to reducing the overall carbon footprint of mobile applications (Goaër, 2021).
Despite the goal for application domain-specific software is still on reducing their energy consumption and carbon footprint, the way to achieve such goals differs from generic software (Figure 8). More precisely, only ML and computer vision are still dealt with high-level sustainability guidelines aimed at reducing their energy consumption and carbon footprint (Figure 8), especially in the model training phase (Brownlee et al., 2021; Pan and Rajan, 2022; Fu et al., 2021; Patterson et al., 2022; Wei et al., 2023). For the other considered application domains, instead, the heat goes over software improvement techniques (Figure 8a), which exploit the application domain-specific context to provide more concrete solutions aimed at reducing their energy consumption and carbon footprint (Figure 8b). This highlights the potentials of narrowing the scope to specific application domains for providing more concrete techniques to design and develop environmentally sustainable software, hence calling for further investigation along this line.
| Highlights | |
| Most of the existing solutions are high-level guidelines and reference models targeting generic software. Their applicability and extension to specific application domains deserves further investigation. | |
| The currently targeted application domains are those known to be computationally intensive (e.g., ML and blockchains) or largely used (e.g., mobile applications). | |
| Narrowing the scope to specific application domains enables providing more concrete techniques to design and develop carbon-efficient software, calling for further investigation along this line. | |
4.5. When?
Figure 9 shows the frequency of lifecycle stages targeted in the selected studies.
Being the present SLR targeting the design and development of environmentally sustainable software, such stages naturally result to be the most targeted by the selected studies. The other three stages are anyhow quite frequent in the selected studies, with requirement analysis targeted by 40 studies, integration & testing targeted by 34 studies, and rollout targeted by 33 studies. This is mainly due to the fact that the currently available reference models and measurement solutions – which together account for more than half of the selected studies – are designed to be applicable across all lifecycle stages, thereby naturally extending their relevance beyond design and development.
In contrast, sustainability guidelines and software improvement techniques tend to be more concentrated on these two stages, with less than half addressing other phases of the lifecycle (Figure 10).
On another front, it is interesting to observe who is involved in which lifecycle stage, according to the state of the art on the design and development of environmentally sustainable software. Figure 10b provides such information, confirming what we observed in Section 4.1: the most involved stakeholders are software developers. Notably, software developers are not only involved in the development stage, but rather result the most involved across all the stages of a software’s lifecycle – which is again in line with what we observed in in Section 4.1, namely that the burden of environmental sustainability is expected to be put on software developers. This, along with the fact that the spanning across all stages seems to apply to all stakeholders, suggests that all stakeholders should be involved – and supported – during all lifecycle stages. Therefore, further investigation should be pursued along this line.
| Highlights | |
| All lifecycle stages result to be significantly covered, despite our focus is on design and development. | |
| Reference models and measurement solutions span across all lifecycle stages, whereas sustainability guidelines and software improvement solutions focus more on design and development only. | |
| Software developers are the most involved in all stages, but all stakeholders span across all lifecycle stages, calling for supporting all stakeholders during all lifecycle stages. | |
4.6. How?
The how question aims at identifying the distribution of publications on environmentally sustainable software design and development, based on bibliometrics like the publication type and year (Figure 11).
After a pioneering journal article in 2001, no activity is reported up to 2009 included. Instead, since 2010 onwards, there were publications on the topic each year, witnessing a renewed interest in the topic, with a peak in 2013 and an ongoing growth since 2020. Overall, conference papers were the most used means of publication. This may be motivated by the authors wishing to timely publish their results, therefore going for conferences that offer a faster publication. An anomaly in this respect is 2022, with more journal articles than conference papers, which may be motivated by the faster publication process recently adopted by international journals.
Notably, the peak in 2013 and the ongoing growth seem to relate two co-occurring events in time. More precisely, the peak in 2013 may be a consequence of a report (OECD, 2009) by OECD (Organisation for Economic Cooperation and Development), which assessed government programmes and business initiatives on ICT and environmental sustainability across 22 OECD countries. Instead, the ongoing growth since 2020 may be a consequence of the recent interest in environmental sustainability. These were raised, e.g., by international organisations like Fridays for Future, and investment policies like the European strategic agenda (European Union, 2019b), which calls for and finances a sustainable growth in European industries, including the IT industry (European Union, 2019a, 2021, 2024).
| Highlights | |
| Conference papers are steadily higher in number than journal articles, mainly because of conferences’ faster publication tracks. | |
| The peak of publication in 2013 may be motivated by the 2009 OECD assessment of government programmes and business initiatives on ICT and environmental sustainability. | |
| An ever increasing number of publications since 2020 witnesses the hotness of the topic, also pairing with the ever increasing interest in environmental sustainability. | |
5. Threats to Validity
Based on the taxonomy developed by Wohlin et al. (Wohlin et al., 2000), four potential threats may apply to the validity of our study. These are the threats to external, internal, construct, and conclusions validities, which we discuss hereafter.
Threats to External Validity. The potential threats to the external validity of a study affect the applicability of the study’s results in a broader and more general context (Wohlin et al., 2000). In our case, since the selected studies are obtained from a large extent of online available sources, our results and observations may be only partly applicable to the broad area of the design and development of carbon efficient software, and this may potentially threaten the external validity of our study. To mitigate this potential threat, we organised multiple feedback sessions during our study. We analysed the discussion following up from each feedback session, and we exploited the resulting qualitative data to fine-tune both our research methods and the applicability of our findings. We also prepared a replication package bundling the intermediary artefacts and the final results produced while running our study, which is publicly available online (Danushi et al., 2025). The replication package, together with the references to the selected studies listed in Table 2, can be used to deepen the understanding of our data, as well as to repeat our study and verify our findings.
Additionally, the external validity of our study may be threatened by the risk of having missed relevant studies, since the concepts related to those included in our search strings are differently named in such studies. To mitigate this potential threat, we have explicitly included all relevant synonyms in our search strings and we also enacted backward snowballing (see Section 2.2).
Threats to Internal Validity The internal validity of a study concerns the validity of the method employed to analyse the study data (Wohlin et al., 2000). To reinforce the internal validity of our study, we relied on established guidelines to conduct systematic literature reviews in software engineering (Kitchenham and Charters, 2007; Petersen et al., 2008, 2015). Additionally, we enacted various triangulation rounds and adopted inter-rater reliability assessment (see Section 2), which were conceived to avoid possible biases by triangulation.
Threats to Construct and Conclusions Validity The construct and conclusions validities concerns the generalisability of the constructs under study and the degree to which the conclusions of a study are reasonably based on the available data, respectively (Wohlin et al., 2000). The construct and conclusions validities of our study may have been affected by observer and interpretation biases. To mitigate these potential threats, we exploited inter-rater reliability assessment (see Section 2), and we also independently drawn conclusions from the available data. Such conclusions were then suitably merged and double-checked against the selected (and related) studies in a joint discussion session.
6. Related Work
Several secondary studies survey the state of the art for sustainable software. A first example is given by Ormazabal et al (Ormazabal et al., 2014), who survey the software tools that can be used to assess a product’s lifecycle and analyse its carbon footprint. Sinha and Chandel (Sinha and Chandel, 2014) and Sudarmadi and Garniwa (Sudarmadi and Garniwa, 2023) review the use of software tools for planning of hybrid renewable energy systems. Popescu et al. (Popescu et al., 2022) survey the existing approaches for building digital twins of manufacturing production systems by incorporating environmental sustainability aspects. The above listed studies differ from ours in their objective, as they discuss how to achieve environmental sustainability with software. We instead focus on the solutions improving the environmental sustainability of software itself.
Other noteworthy secondary studies are the literature reviews by Salam and Khan (Salam and Khan, 2016) and Sambhanthan and Potdar (Sambhanthan and Potdar, 2016). Salam and Khan (Salam and Khan, 2016) classify and analyse the success factors for vendors of green software. Sambhanthan and Potdar (Sambhanthan and Potdar, 2016) instead focus on the existing frameworks for assessing a business’ sustainability, therein included ICT sustainability. Therefore, the reviews by Salam and Khan (Salam and Khan, 2016) and Sambhanthan and Potdar (Sambhanthan and Potdar, 2016) differ from ours in the objective, as they focus on the business-related aspects of software’s sustainability. We instead aim at classifying the technical solutions for designing and developing carbon-efficient software.
Andrikopoulos et al. (Andrikopoulos et al., 2022) provide a systematic mapping study aimed at analysing which architecting activities are affected/affect which of the four different dimensions of sustainability (i.e., technical, economic, environmental, and social). Ahmadisakha and Andrikopoulos (Ahmadisakha and Andrikopoulos, 2024) deepen into the topic, focussing specifically on Cloud-based software architecture, still analysing how architectural decisions for Cloud software are affected/affect the four different dimensions of sustainability. Therefore, the scope and goal of both the above secondary studies differ from ours, as they aim at eliciting which of the four different dimensions of sustainability impact on/are impacted by which architecting phase. We instead focus on the 5W1H of designing and developing carbon-efficient software. Additionally, Ahmadisakha and Andrikopoulos (Ahmadisakha and Andrikopoulos, 2024) focus on Cloud-based software, while we rather consider multiple different application domains.
Other secondary studies focus on improving the environmental sustainability of software in specific target application domains. For example, Azimzadeh and Tabrizi (Azimzadeh and Tabrizi, 2015) and Khan et al. (Khan et al., 2020) provide two different reviews on software techniques for managing green Cloud data centres. Gaglianese et al. (Gaglianese et al., 2023b) survey the existing approaches for the environmentally sustainable orchestration of containerised software applications over Cloud-Edge infrastructures. Castaño et al. (Castaño et al., 2023) analyse and classify the carbon footprint of training publicly available ML models. Verdecchia et al. (Verdecchia et al., 2023) review the state of the art on green AI, by distinguishing the proposals supporting the training and operation of AI models. López-Pérez et al. (López-Pérez et al., 2022) survey the existing techniques for operating 5G networks in an energy-efficient manner. Thus, all the above-mentioned secondary studies focus on a specific application domain and on different lifecycle stages than ours, as we rather consider designing and developing carbon-efficient software by covering – and classifying – multiple different application domains.
The SLRs by Bozzelli et al. (Bozzelli et al., 2014), Calero et al. (Calero et al., 2013), Koziolek(Koziolek, 2011), and Venters et al. (Venters et al., 2018, 2023) are a step closer to ours, as they all pertain to the broader scopus of environmental sustainability in software engineering. More precisely, Bozzelli et al. (Bozzelli et al., 2014) and Koziolek (Koziolek, 2011) classify the existing metrics for estimating the greenness of software. Calero et al. (Calero et al., 2013) elicit the software quality aspects related to environmental sustainability. Venters et al. (Venters et al., 2018, 2023) go a step further, by also identifying which definitions of software sustainability are available, as well as how software sustainability impacts on architectural decisions, goals, and education. Thus, even if slightly overlapping, the objectives of the above mentioned SLRs differ from ours, as we rather aim at answering to the 5W1H questions for designing and developing carbon-efficient software (see Section 2.1).
Similar arguments apply to the SLR by Penzenstadler et al. (Penzenstadler et al., 2012), who started with similar aims than ours, to then realise that – at the time when the study was conducted – only a few studies tackled environmental sustainability in software engineering. The SLR by Penzenstadler was then redirected to identify and classify the broader scope of relating environmental sustainability to software, one of these being how it relates with software engineering. Therefore, the SLR by Penzenstadler et al. (Penzenstadler et al., 2012) differs from ours in the ultimate contribution, as – after more than 10 years – the literature on sustainability is more mature, as we showed by focusing on the narrower scope of designing and developing carbon-efficient software.
In summary, to the best of our knowledge, ours is the first SLR on the state of the art on designing and developing carbon-efficient software. It is the first doing it considering not only the what of sustainable software design and development, but rather aiming at answering to the all 5W1H questions related to designing and developing carbon-efficient software (see Section 2.1). This enabled us, e.g., to first reason on the stakeholders for carbon-efficient software design and development, to span over multiple application domains, and to first distinguish whether environmental sustainability is dealt with by only reducing energy consumption or by also considering the carbon footprint of software.
7. Conclusions and Outlook
We presented the results of a SLR on the design and development of carbon-efficient software. After classifying and recapping the existing solutions for the tasks, we analysed who are the targeted stakeholders, what methods allow to design and develop carbon-efficient software, why, where and when the proposed solutions are applied (in terms of application domains and lifecycle stages, respectively), and how the selected studies have been presented to the scientific community.
The results of our SLR can help researchers and practitioners identify and adopt existing guidelines and reference models for designing and developing carbon-efficient software, as well as to select applicable solutions for their software projects. Our analysis also highlights a growing interest in this topic since 2020, providing a basis for establishing future research directions. To facilitate this, we outline seven research directions derived from our analysis, each mapped to the highlights in Section 4 that motivate these directions.
Relieving the burden from software developers (–). The existing solutions mainly aim at supporting software developers, and this somehow puts the burden of designing and developing carbon-efficient software on software developers themselves. There is a need for solutions to support and involve all the potential stakeholders in the process. In particular, we should enable software architects and sustainability engineers to get more involved in the design of carbon-efficient software, taking responsibility for sustainability-oriented design choices, thus complementing the work and efforts carried out by software developers.
Standardising carbon efficiency for software (). Most of the surveyed proposals are reference models, witnessing the need for clarifying the notion of carbon efficiency for software, also outlining how this can be used to support the design and development of more sustainable software. However, despite sharing some base concepts, the currently available reference models differ one from another, calling for standardisation efforts. The availability of a standardised reference model for software carbon efficiency would enable all stakeholders to rely on the same understanding of carbon efficiency. It would also lay the foundations for devising full-fledged solutions to support the design and development of carbon-efficient software.
Finding a trade-off between applicability and carbon efficiency (). Most of the existing solutions for designing and developing carbon-efficient software adopt a black-box approach. This is mainly motivated by applicability reasons, as black-box approaches can work even with software whose sources are not available, e.g., for privacy or legacy reasons. At the same time, while surveying the state of the art, we observed some potentials of white-box analyses. For example, Le Goaër et al (Goaër, 2021, 2023; Goaër and Hertout, 2022) highlighted the potential of defining, detecting, and resolving so-called green code smells, namely possible symptoms of bad coding decisions that may worsen the environmental impact of software. Mehra et al. (Mehra et al., 2023) instead showcased how static analysis enabled them to successfully resolve code inefficiencies in an industrial use case. Therefore, the potentials of white-box approaches deserve further investigation, in an effort to find a suitable trade-off between the applicability of a solution and its effectiveness in supporting the design and development of carbon-efficient software.
Raising the overall TRL (). We observed that, overall, the TRL of the state of the art on carbon-efficient software design and development is relatively low (with only one out of 65 selected studies reaching a TRL 5). To favour the adoption of existing – and newly proposed – solutions by the software industry, the overall TRL should be increased. In particular, the existing proposals should be prototyped and assessed over industry-relevant scenarios, to then release more mature tooling to the software industry. To this end, there is also a need for benchmarks, industry-relevant use cases, and realistic case studies that could be used to thoroughly validate existing/newly proposed solutions for designing and developing carbon-efficient software.
Increasing carbon-awareness (). Most of the existing solutions consider energy consumption as the main discriminant to improve the environmental sustainability of software. This is motivated by the shared assumption that, by directly reducing the energy consumption of software, its carbon footprint gets indirectly reduced as well. However, energy mixes play a crucial role to relate energy consumption with carbon footprint, as different energy mixes come with different associated carbon intensities (Aiello et al., 2024). This may impair the effectiveness of directly considering energy consumption to indirectly reduce carbon emission, since e.g., when software is powered with clean and renewable energy, reducing its energy consumption does not significantly change its carbon footprint (Forti et al., 2024). This calls for further research along the line of directly considering energy mixes and carbon emissions of software, rather than only focussing on reducing its carbon footprint by limiting its energy consumption.
Adopting existing solutions in application domains (,–). Most of the existing solutions are high-level guidelines, reference models, and measurement solutions for generic software. Being them generic, their concrete adoption in specific application domains may require considerable efforts, e.g., to adapt them to work with the peculiar aspects that mostly impact on the carbon footprint on application domain-specific software (like the training phase for ML models, or the PUE for Cloud data centers). This calls for identifying and considering such peculiar aspects, to define/improve domain-specific guidelines and reference models, as well as to adapt existing techniques to work with application domain-specific software. This would enable the effective use and assessment of existing solutions available in multiple different application domains, thus possibly paving the way towards their adoption. Additionally, narrowing the scope to specific application domains showed the potential to not only reduce, but also minimise, the carbon efficiency of the software under design/development.
Relating design and development with all lifecycle stages (–). Even if focussing on the design and development of carbon-efficient software, this SLR highlighted how the existing solutions span across all the stages of the software lifecycle. This is motivated by the fact that sustainability guidelines and reference models are often proposed for the full software lifecycle. Going beyond this, we observed the effectiveness of the design and development stages in achieving carbon-efficient impacts on/is impacted by the other lifecycle stages as well. This calls for supporting the software design and development stages by considering, keeping track of, and assessing their impacts/relations with the whole software lifecycle.
Additionally, although this survey focuses specifically on the design and development stages, the outcomes discussed in Section 4.5 highlight that other lifecycle stages might contribute to the carbon-efficiency of software artifacts, e.g., requirements engineering might set carbon-efficiency goals, testing might help validating carbon-efficiency, and operation might further reduce the carbon footprint of running software. While the carbon-efficient operation of software has already been analysed, e.g., in Gaglianese et al. (Gaglianese et al., 2023b), future secondary studies could complement the results of this work by exploring how the other lifecycle stages can further contribute to the carbon-efficiency of software artifacts. Such secondary studies are left in the scope of future work.
Acknowledgements.
This work has been partly supported by projects: FREEDA (CUP: I53D23003550006), funded by the frameworks PRIN (MUR, Italy) and Next Generation EU and Energy-aware management of software applications in Cloud-IoT ecosystems (RIC2021_PON_A18) funded over ESF REACT-EU resources by the Italian Ministry of University and Research through PON Ricerca e Innovazione 2014–20.References
- (1)
- Adisa et al. (2024) Mikhail O. Adisa, Shola Oyedeji, and Jari Porras. 2024. The nexus between ICT, top-down and bottom-up approaches for sustainability activities: A systematic mapping study. Journal of Cleaner Production 449 (2024), 141768. https://doi.org/10.1016/j.jclepro.2024.141768
- Ahmadisakha and Andrikopoulos (2024) Sahar Ahmadisakha and Vasilios Andrikopoulos. 2024. Architecting for sustainability of and in the cloud: A systematic literature review. Inf. Softw. Technol. 171 (2024), 107459. https://doi.org/10.1016/J.INFSOF.2024.107459
- Ahmed (2020) Shakeel Ahmed. 2020. Environmental sustainability coding techniques for cloud computing. Int. Journal of Advanced Computer Science and Applications 11, 10 (2020), 231 – 237. https://doi.org/10.14569/IJACSA.2020.0111031
- Aiello et al. (2024) Marco Aiello, Tobias Piontek, and Kawsar Haghshenas. 2024. Carbon emission-aware job scheduling for Kubernetes deployments. The Journal of Supercomputing 80 (2024). https://doi.org/10.1007/s11227-023-05506-7
- Alofi et al. (2023) Akram Alofi, Mahmoud A. Bokhari, Rami Bahsoon, and Robert Hendley. 2023. Self-Optimizing the Environmental Sustainability of Blockchain-Based Systems. IEEE Transactions on Sustainable Computing (2023), 1–13. https://doi.org/10.1109/TSUSC.2023.3325881 In press.
- Alofi et al. (2022) Akram Alofi, Mahmoud A. Bokhari, Rami Bahsoon, and Robert J. Hendley. 2022. Optimizing the Energy Consumption of Blockchain-Based Systems Using Evolutionary Algorithms: A New Problem Formulation. IEEE Transactions on Sustainable Computing 7, 4 (2022), 910–922. https://doi.org/10.1109/TSUSC.2022.3160491
- Anderson et al. (2023) Thomas Anderson, Adam Belay, Mosharaf Chowdhury, Asaf Cidon, and Irene Zhang. 2023. Treehouse: A case for carbon-aware datacenter software. ACM SIGENERGY Energy Informatics Review 3, 3 (2023), 64–70.
- Andrikopoulos et al. (2022) Vasilios Andrikopoulos, Rares-Dorian Boza, Carlos Perales, and Patricia Lago. 2022. Sustainability in Software Architecture: A Systematic Mapping Study. In 48th Euromicro Conference on Software Engineering and Advanced Applications, SEAA 2022, Maspalomas, Gran Canaria, Spain, 31 August - 2 September 2022, Gustavo Marrero Callicó et al. (Eds.). IEEE, 426–433. https://doi.org/10.1109/SEAA56994.2022.00073
- Atkinson et al. (2014) Colin Atkinson, Thomas Schulze, and Sonja Klingert. 2014. Facilitating Greener IT through Green Specifications. IEEE Software 31, 3 (2014), 56–63. https://doi.org/10.1109/MS.2014.19
- Azad and Hyrynsalmi (2023) Nasreen Azad and Sami Hyrynsalmi. 2023. DevOps critical success factors — A systematic literature review. Information and Software Technology 157 (2023), 107150. https://doi.org/10.1016/j.infsof.2023.107150
- Azimzadeh and Tabrizi (2015) Aryan Azimzadeh and Nasseh Tabrizi. 2015. A Taxonomy and Survey of Green Data Centers. In 2015 International Conf. on Computational Science and Computational Intelligence (CSCI). 128–131. https://doi.org/10.1109/CSCI.2015.70
- Becker et al. (2016) Christoph Becker, Stefanie Betz, Ruzanna Chitchyan, Leticia Duboc, Steve M. Easterbrook, Birgit Penzenstadler, Norbert Seyff, and Colin C. Venters. 2016. Requirements: The Key to Sustainability. IEEE Software 33, 1 (2016), 56–65. https://doi.org/10.1109/MS.2015.158
- Beghoura et al. (2014) Mohamed Amine Beghoura, Abdelhak Boubetra, and Abdallah Boukerram. 2014. Green Applications Awareness: NonLinear Energy Consumption Model for Green Evaluation. In Eighth Int. Conf. on Next Generation Mobile Apps, Services and Technologies, NGMAST 2014, Oxford, United Kingdom, September 10-12, 2014. IEEE, 48–53. https://doi.org/10.1109/NGMAST.2014.33
- Belkhir and Elmeligi (2018) Lotfi Belkhir and Ahmed Elmeligi. 2018. Assessing ICT global emissions footprint: Trends to 2040 & recommendations. Journal of Cleaner Production 177 (2018), 448–463. https://doi.org/10.1016/j.jclepro.2017.12.239
- Bozzelli et al. (2014) Paolo Bozzelli, Qing Gu, and Lago Patricia. 2014. A systematic literature review on green software metrics. VU Technial Report, https://research.vu.nl/ws/portalfiles/portal/910331.
- Brownlee et al. (2021) Alexander E. I. Brownlee, Jason Adair, Saemundur O. Haraldsson, and John Jabbo. 2021. Exploring the Accuracy - Energy Trade-off in Machine Learning. In 10th IEEE/ACM Int. Workshop on Genetic Improvement, GI@ICSE 2021, Madrid, Spain, May 30, 2021. IEEE, 11–18. https://doi.org/10.1109/GI52543.2021.00011
- Brundtland et al. (1987) Gro Harlem Brundtland, Mansour Khalid, et al. 1987. Our common future. Oxford University Press, Oxford, GB.
- Caballar (2024) Rina Diane Caballar. 2024. We Need to Decarbonize Software: The Way We Write Software has Unappreciated Environmental Impacts. IEEE Spectrum 61, 4 (2024), 26–31. https://doi.org/10.1109/MSPEC.2024.10491388
- Calero et al. (2013) Coral Calero, Manuel F. Bertoa, and María Ángeles Moraga. 2013. A systematic literature review for software sustainability measures. In 2nd International Workshop on Green and Sustainable Software, GREENS 2013, San Francisco, CA, USA, May 20, 2013, Patricia Lago et al. (Eds.). IEEE Computer Society, 46–53. https://doi.org/10.1109/GREENS.2013.6606421
- Castaño et al. (2023) Joel Castaño, Silverio Martínez-Fernández, Xavier Franch, and Justus Bogner. 2023. Exploring the Carbon Footprint of Hugging Face’s ML Models: A Repository Mining Study. In ACM/IEEE International Symposium on Empirical Software Engineering and Measurement, ESEM 2023, New Orleans, LA, USA, October 26-27, 2023. IEEE, 1–12. https://doi.org/10.1109/ESEM56168.2023.10304801
- Chauhan and Saxena (2013) Nitin Singh Chauhan and Ashutosh Saxena. 2013. A Green Software Development Life Cycle for Cloud Computing. IT Professional 15, 1 (2013), 28–34. https://doi.org/10.1109/MITP.2013.6
- Condori-Fernández and Lago (2018) Nelly Condori-Fernández and Patricia Lago. 2018. The Influence of Green Strategies Design onto Quality Requirements Prioritization. In Requirements Engineering: Foundation for Software Quality - 24th Int. Working Conf., REFSQ 2018, Utrecht, The Netherlands, March 19-22, 2018, Proceedings (Lecture Notes in Computer Science, Vol. 10753), Erik Kamsties et al. (Eds.). Springer, 189–205. https://doi.org/10.1007/978-3-319-77243-1_12
- Danushi et al. (2025) Ornela Danushi, Stefano Forti, and Jacopo Soldani. 2025. Carbon-Efficient Software Design and Development: Replication Package. Zenodo, https://doi.org/10.5281/zenodo.14917055.
- Dennard et al. (1974) Robert H. Dennard et al. 1974. Design Of Ion-implanted MOSFET’s with Very Small Physical Dimensions. IEEE Journal of Solid-State Circuits 9, 5 (1974), 256 – 268. https://doi.org/10.1109/JSSC.1974.1050511
- Dick et al. (2013) Markus Dick, Jakob Drangmeister, Eva Kern, and Stefan Naumann. 2013. Green software engineering with agile methods. In 2nd Int. Workshop on Green and Sustainable Software, GREENS 2013, San Francisco, CA, USA, May 20, 2013, Patricia Lago et al. (Eds.). IEEE Computer Society, 78–85. https://doi.org/10.1109/GREENS.2013.6606425
- Dick et al. (2010) Markus Dick, Stefan Naumann, and Norbert Kuhn. 2010. A Model and Selected Instances of Green and Sustainable Software. In What Kind of Information Society? Governance, Virtuality, Surveillance, Sustainability, Resilience - 9th IFIP TC 9 Int. Conf., HCC9 2010 and 1st IFIP TC 11 Int. Conf., CIP 2010, Held as Part of WCC 2010, Brisbane, Australia, September 20-23, 2010. Proceedings (IFIP Advances in Information and Communication Technology, Vol. 328), Jacques Berleur et al. (Eds.). Springer, 248–259. https://doi.org/10.1007/978-3-642-15479-9_24
- European Commission (2014) European Commission. 2014. Technology readiness levels (TRL). https://ec.europa.eu/research/participants/data/ref/h2020/wp/2014_2015/annexes/h2020-wp1415-annex-g-trl_en.pdf.
- European Union (2019a) European Union. 2019a. The European Green Deal. https://commission.europa.eu/strategy-and-policy/priorities-2019-2024/european-green-deal_en.
- European Union (2019b) European Union. 2019b. Strategic Agenda for 2019-2024. https://www.consilium.europa.eu/en/eu-strategic-agenda-2019-2024/.
- European Union (2021) European Union. 2021. Financing the Climate Transition. https://www.consilium.europa.eu/en/policies/climate-finance/.
- European Union (2024) European Union. 2024. EU industrial policy. https://www.consilium.europa.eu/en/policies/eu-industrial-policy/.
- Fakhar et al. (2012) Faiza Fakhar, Barkha Javed, Raihan Ur Rasool, Owais Ahmed Malik, and Khurram Zulfiqar. 2012. Software level green computing for large scale systems. Journal of Cloud Computing 1 (2012), 4. https://doi.org/10.1186/2192-113X-1-4
- Forti et al. (2024) Stefano Forti, Jacopo Soldani, and Antonio Brogi. 2024. Carbon-aware Software Services. arXiv:2405.12582 [cs.SE] https://arxiv.org/abs/2405.12582
- Fu et al. (2021) Andre Fu, Mahdi S. Hosseini, and Konstantinos N. Plataniotis. 2021. Reconsidering CO2 Emissions From Computer Vision. In IEEE Conf. on Computer Vision and Pattern Recognition Workshops, CVPR Workshops 2021, virtual, June 19-25, 2021. Computer Vision Foundation / IEEE, 2311–2317. https://doi.org/10.1109/CVPRW53098.2021.00262
- Gaglianese et al. (2023a) Marco Gaglianese et al. 2023a. Green Orchestration of Cloud-Edge Applications: State of the Art and Open Challenges. In IEEE International Conference on Service-Oriented System Engineering, SOSE 2023. IEEE, 250–261. https://doi.org/10.1109/SOSE58276.2023.00036
- Gaglianese et al. (2023b) Marco Gaglianese, Jacopo Soldani, Stefano Forti, and Antonio Brogi. 2023b. Green Orchestration of Cloud-Edge Applications: State of the Art and Open Challenges. In IEEE International Conference on Service-Oriented System Engineering, SOSE 2023, Athens, Greece, July 17-20, 2023. IEEE, 250–261. https://doi.org/10.1109/SOSE58276.2023.00036
- Goaër (2021) Olivier Le Goaër. 2021. Enforcing green code with Android lint. In Proceedings of the 35th IEEE/ACM Int. Conf. on Automated Software Engineering. Association for Computing Machinery, New York, NY, USA, 85–90. https://doi.org/10.1145/3417113.3422188
- Goaër (2023) Olivier Le Goaër. 2023. Decarbonizing Software with Free and Open Source Software: The ecoCode Project. In Proceedings of the Research Projects Exhibition Papers Presented at the 35th Int. Conf. on Advanced Information Systems Engineering (CAiSE 2023), Zaragoza, Spain, June 12-16, 2023 (CEUR Workshop Proceedings, Vol. 3413), Jaime Font et al. (Eds.). CEUR-WS.org, 9–13. https://ceur-ws.org/Vol-3413/paper2.pdf
- Goaër and Hertout (2022) Olivier Le Goaër and Julien Hertout. 2022. ecoCode: a SonarQube Plugin to Remove Energy Smells from Android Projects. In 37th IEEE/ACM Int. Conf. on Automated Software Engineering, ASE 2022, Rochester, MI, USA, October 10-14, 2022. ACM, 157:1–157:4. https://doi.org/10.1145/3551349.3559518
- Heithoff et al. (2023) Malte Heithoff, Alexander David Hellwig, Judith Michael, and Bernhard Rumpe. 2023. Digital Twins for Sustainable Software Systems. In 7th IEEE/ACM Int. Workshop on Green And Sustainable Software, GREENS@ICSE 2023, Melbourne, Australia, May 14, 2023. IEEE, 19–23. https://doi.org/10.1109/GREENS59328.2023.00010
- Hoesch-Klohe et al. (2010) Konstantin Hoesch-Klohe, Aditya Ghose, and Lam-Son Lê. 2010. Towards Green Business Process Management. In 2010 IEEE Int. Conf. on Services Computing, SCC 2010, Miami, Florida, USA, July 5-10, 2010. IEEE Computer Society, 386–393. https://doi.org/10.1109/SCC.2010.21
- Ibrahim et al. (2021) Siti Rohana Ahmad Ibrahim, Jamaiah Yahaya, Hasimi Salehudin, and Aziz Deraman. 2021. The Development of Green Software Process Model: A Qualitative Design and Pilot Study. Int. Journal of Advanced Computer Science and Applications 12, 8 (2021), 589 – 598. https://doi.org/10.14569/IJACSA.2021.0120869
- Ibrahim et al. (2022) Siti Rohana Ahmad Ibrahim, Jamaiah Yahaya, and Hasimi Sallehudin. 2022. Green Software Process Factors: A Qualitative Study. Sustainability 14, 18 (2022). https://doi.org/10.3390/su141811180
- Jayanthi et al. (2021) Harika Jayanthi, Anjali Kulkarni, Aishwarya E Patil, and Sheela S V. 2021. An Organizational Structure for Sustainable Software Development. In Proceedings - 2021 3rd Int. Conf. on Advances in Computing, Communication Control and Networking, ICAC3N 2021. 1539 – 1542. https://doi.org/10.1109/ICAC3N53548.2021.9725722
- Jebraeil et al. (2017) Nigar Jebraeil, Adel Noureddine, Joseph Doyle, Syed Islam, and Rabih Bashroush. 2017. gUML: Reasoning about Energy at Design Time by Extending UML Deployment Diagrams with Data Centre Contextual Information. In 2017 IEEE World Congress on Services, SERVICES 2017, Honolulu, HI, USA, June 25-30, 2017, Rami Bahsoon and Zhixiong Chen (Eds.). IEEE Computer Society, 59–66. https://doi.org/10.1109/SERVICES.2017.19
- Johann et al. (2011) Timo Johann, Markus Dick, Eva Kern, and Stefan Naumann. 2011. Sustainable development, sustainable software, and sustainable software engineering: An integrated approach. In 2011 Int. Symposium on Humanities, Science and Engineering Research. 34–39. https://doi.org/10.1109/SHUSER.2011.6008495
- Karunakaran and Rao (2013) Deepak Karunakaran and G Subrahmanya Vrk Rao. 2013. A petri net simulation of software development lifecycle towards green IT. In 2013 IEEE Conf. on Open Systems, ICOS 2013. 58 – 62. https://doi.org/10.1109/ICOS.2013.6735048
- Kern et al. (2013) Eva Kern, Markus Dick, Stefan Naumann, Achim Guldner, and Timo Johann. 2013. Green Software and Green Software Engineering – Definitions, Measurements, and Quality Aspects. https://api.semanticscholar.org/CorpusID:40610952
- Kern et al. (2015) Eva Kern, Markus Dick, Stefan Naumann, and Tim Hiller. 2015. Impacts of software and its engineering on the carbon footprint of ICT. Environmental Impact Assessment Review 52 (2015), 53–61. https://doi.org/10.1016/j.eiar.2014.07.003
- Kern et al. (2018) Eva Kern, Lorenz M. Hilty, Achim Guldner, Yuliyan V. Maksimov, Andreas Filler, Jens Gröger, and Stefan Naumann. 2018. Sustainable software products - Towards assessment criteria for resource and energy efficiency. Future Generation Computer Systems 86 (2018), 199–210. https://doi.org/10.1016/J.FUTURE.2018.02.044
- Khan et al. (2020) Fauzia Khan, Hina Anwar, Dietmar Pfahl, and Satish Narayana Srirama. 2020. Software Techniques for Making Cloud Data Centers Energy-efficient: A Systematic Mapping Study. In 46th Euromicro Conference on Software Engineering and Advanced Applications, SEAA 2020, Portoroz, Slovenia, August 26-28, 2020. IEEE, 479–486. https://doi.org/10.1109/SEAA51224.2020.00081
- Kharchenko and Gorbenko (2013) Vyacheslav S. Kharchenko and Anatoliy Gorbenko. 2013. Evolution of von Neumann’s paradigm: Dependable and green computing. In East-West Design & Test Symposium, EWDTS 2013, Rostov-on-Don, Russia, September 27-30, 2013. IEEE Computer Society, 1–6. https://doi.org/10.1109/EWDTS.2013.6673090
- Kipp et al. (2011) Alexander Kipp, Tao Jiang, and Mariagrazia Fugini. 2011. Green Metrics for Energy-aware IT Systems. In Int. Conf. on Complex, Intelligent and Software Intensive Systems, CISIS 2011, June 30 - July 2, 2011, Korean Bible University, Seoul, Korea. IEEE Computer Society, 241–248. https://doi.org/10.1109/CISIS.2011.42
- Kipp et al. (2012) Alexander Kipp, Tao Jiang, Mariagrazia Fugini, and Ioan Salomie. 2012. Layered Green Performance Indicators. Future Generation Computer Systems 28, 2 (2012), 478–489. https://doi.org/10.1016/J.FUTURE.2011.05.005
- Kitchenham and Charters (2007) Barbara Ann Kitchenham and Stuart Charters. 2007. Guidelines for performing Systematic Literature Reviews in Software Engineering. Keele University and Durham University Joint Report, EBSE 2007-001.
- Kneuper (2018) Ralf Kneuper. 2018. Software Processes and Life Cycle Models (1 ed.). Springer Cham. 345 pages. https://doi.org/doi.org/10.1007/978-3-319-98845-0
- Koçak and Alptekin (2019) Sedef Akinli Koçak and Gülfem Isiklar Alptekin. 2019. A Utility Model for Designing Environmentally Sustainable Software. In Proceedings of the 8th Int. Workshop on Requirements Engineering for Sustainable Systems co-located with 27th Int. Conf. on Requirements Engineering, Jeju, South Korea, September 24, 2019 (CEUR Workshop Proceedings, Vol. 2541), Birgit Penzenstadler et al. (Eds.). CEUR-WS.org. https://ceur-ws.org/Vol-2541/paper8.pdf
- Koziolek (2011) Heiko Koziolek. 2011. Sustainability evaluation of software architectures: a systematic review. In 7th International Conference on the Quality of Software Architectures, QoSA 2011 and 2nd International Symposium on Architecting Critical Systems, ISARCS 2011. Boulder, CO, USA, June 20-24, 2011, Proceedings, Ivica Crnkovic et al. (Eds.). ACM, 3–12. https://doi.org/10.1145/2000259.2000263
- Krippendorff (2004) Klaus Krippendorff. 2004. Content Analysis: An Introduction to Its Methodology (2 ed.). Sage, Thousand Oaks, CA.
- Lago (2019) Patricia Lago. 2019. Architecture design decision maps for software sustainability. In Proceedings of the 41st Int. Conf. on Software Engineering: Software Engineering in Society, ICSE 2019, Montreal, QC, Canada, May 25-31, 2019, Rick Kazman and Liliana Pasquale (Eds.). ACM, 61–64. https://doi.org/10.1109/ICSE-SEIS.2019.00015
- Lago and Jansen (2010) Patricia Lago and Toon Jansen. 2010. Creating Environmental Awareness in Service Oriented Software Engineering. In Service-Oriented Computing - ICSOC 2010 Int. Workshops, PAASC, WESOA, SEE, and SOC-LOG, San Francisco, CA, USA, December 7-10, 2010, Revised Selected Papers (Lecture Notes in Computer Science, Vol. 6568), E. Michael Maximilien et al. (Eds.). 181–186. https://doi.org/10.1007/978-3-642-19394-1_19
- Lami and Buglione (2012) Giuseppe Lami and Luigi Buglione. 2012. Measuring Software Sustainability from a Process-centric Perspective. In 2012 Joint Conf. of the 22nd Int. Workshop on Software Measurement and the 2012 Seventh Int. Conf. on Software Process and Product Measurement, Assisi, Italy, October 17-19, 2012. IEEE Computer Society, 53–59. https://doi.org/10.1109/IWSM-MENSURA.2012.16
- Lami et al. (2014) Giuseppe Lami, Fabrizio Fabbrini, and Luigi Buglione. 2014. An ISO/IEC 33000-Compliant Measurement Framework for Software Process Sustainability Assessment. In 2014 Joint Conf. of the Int. Workshop on Software Measurement and the Int. Conf. on Software Process and Product Measurement, Rotterdam, The Netherlands, October 6-8, 2014. IEEE Computer Society, 50–59. https://doi.org/10.1109/IWSM.MENSURA.2014.34
- Lami et al. (2013) Giuseppe Lami, Fabrizio Fabbrini, and Mario Fusani. 2013. A methodology to derive sustainability indicators for software development projects. In Int. Conf. on Software and System Process, ICSSP ’13, San Francisco, CA, USA, May 18-19, 2013, Jürgen Münch et al. (Eds.). ACM, 70–77. https://doi.org/10.1145/2486046.2486060
- Lasla et al. (2022) Noureddine Lasla, Lina Al-Sahan, Mohamed Abdallah, and Mohamed Younis. 2022. Green-PoW: An energy-efficient blockchain Proof-of-Work consensus algorithm. Comput. Netw. 214, C (2022). https://doi.org/10.1016/j.comnet.2022.109118
- López-Pérez et al. (2022) David López-Pérez, Antonio De Domenico, Nicola Piovesan, Xinli Geng, Harvey Bao, Qitao Song, and Mérouane Debbah. 2022. A Survey on 5G Radio Access Network Energy Efficiency: Massive MIMO, Lean Carrier Design, Sleep Modes, and Machine Learning. IEEE Commun. Surv. Tutorials 24, 1 (2022), 653–697. https://doi.org/10.1109/COMST.2022.3142532
- Mahmoud and Ahmad (2013) Sara Mahmoud and Imtiaz Ahmad. 2013. A Green Model for Sustainable Software Engineering. Int. Journal of Software Engineering and Its Applications 7, 4 (2013).
- Mancebo et al. (2021) Javier Mancebo, Coral Calero, Félix García, María Ángeles Moraga, and Ignacio García Rodríguez de Guzmán. 2021. FEETINGS: Framework for Energy Efficiency Testing to Improve eNvironmental Goals of the Software. In Software Sustainability, Coral Calero et al. (Eds.). Springer, 69–87. https://doi.org/10.1007/978-3-030-69970-3_4
- Manner (2022) Jukka Manner. 2022. Black software — the energy unsustainability of software systems in the 21st century. Oxford Open Energy 2 (12 2022), oiac011. https://doi.org/10.1093/ooenergy/oiac011
- Mehra et al. (2023) Rohit Mehra, Priyavanshi Pathania, Vibhu Saujanya Sharma, Vikrant Kaulgud, Sanjay Podder, and Adam P. Burden. 2023. Assessing the Impact of Refactoring Energy-Inefficient Code Patterns on Software Sustainability: An Industry Case Study. In 38th IEEE/ACM Int. Conf. on Automated Software Engineering, ASE 2023, Luxembourg, September 11-15, 2023. IEEE, 1825–1827. https://doi.org/10.1109/ASE56229.2023.00205
- Mehra et al. (2022) Rohit Mehra, Vibhu Saujanya Sharma, Vikrant Kaulgud, Sanjay Podder, and Adam P. Burden. 2022. Towards a Green Quotient for Software Projects. In 44th IEEE/ACM Int. Conf. on Software Engineering: Software Engineering in Practice, ICSE (SEIP) 2022, Pittsburgh, PA, USA, May 22-24, 2022. IEEE, 295–296. https://doi.org/10.1109/ICSE-SEIP55303.2022.9793921
- Mingay and Pamlin (2008) Simon Mingay and Dennis Pamlin. 2008. Assessment of Global Low-Carbon and Environmental Leadership in the ICT Sector. Vol. ID G00161337, Gartner and WWF.
- Mohankumar and Kumar (2016) M. Mohankumar and Anand M. Kumar. 2016. Green based Software Development Life Cycle Model for Software Engineering. Indian Journal of Science and Technology 9, 32 (2016), 1–8. https://doi.org/10.17485/ijst/2016/v9i32/88499
- Möller et al. (2024) Tessa Möller, Annika Ernest Högner, Carl-Friedrich Schleussner, Samuel Bien, Niklas H. Kitzmann, Robin D. Lamboll, Joeri Rogelj, Jonathan F. Donges, Johan Rockström, and Nico Wunderling. 2024. Achieving net zero greenhouse gas emissions critical to limit climate tipping risks. Nature Communications 15 (2024). Issue 1. https://doi.org/10.1038/s41467-024-49863-0
- Moore (1998) Gordon E. Moore. 1998. Cramming More Components Onto Integrated Circuits. Proc. IEEE 86, 1 (1998), 82–85. https://doi.org/10.1109/JPROC.1998.658762
- Moshnyaga (2013a) Vasily G. Moshnyaga. 2013a. An assessment of software lifecycle energy. In 2013 23rd Int. Workshop on Power and Timing Modeling, Optimization and Simulation (PATMOS), Karlsruhe, Germany, September 9-11, 2013. IEEE, 112–119. https://doi.org/10.1109/PATMOS.2013.6662163
- Moshnyaga (2013b) Vasily G. Moshnyaga. 2013b. Assessment of Software Lifecycle Energy and its Contribution to Green House Gas Emissions. In 2013 Int. Conf. on Cloud and Green Computing, Karlsruhe, Germany, September 30 - October 2, 2013. IEEE Computer Society, 197–198. https://doi.org/10.1109/CGC.2013.37
- Moshnyaga (2013c) Vasily G. Moshnyaga. 2013c. Lifecycle energy assessment of mobile applications. In Proceedings of the 2013 Int. Workshop on Software Development Lifecycle for Mobile, DeMobile 2013, Saint Petersburg, Russia, August 19, 2013, Aharon Abadi et al. (Eds.). ACM, 17–23. https://doi.org/10.1145/2501553.2501557
- Muralidhar et al. (2022) Rajeev Muralidhar, Renata Borovica-Gajic, and Rajkumar Buyya. 2022. Energy Efficient Computing Systems: Architectures, Abstractions and Modeling to Techniques and Standards. ACM Comput. Surv. 54, 11s (2022), 236:1–236:37. https://doi.org/10.1145/3511094
- Murugesan (2021) San Murugesan. 2021. 10 Recommendations for Green Software Development. Green Software Foundation, https://greensoftware.foundation/articles/10-recommendations-for-green-software-development.
- Naumann et al. (2011) Stefan Naumann, Markus Dick, Eva Kern, and Timo Johann. 2011. The GREENSOFT Model: A reference model for green and sustainable software and its engineering. Sustainable Computing: Informatics and Systems 1, 4 (2011), 294–304. https://doi.org/10.1016/J.SUSCOM.2011.06.004
- Naumann et al. (2015) Stefan Naumann, Eva Kern, Markus Dick, and Timo Johann. 2015. Sustainable Software Engineering: Process and Quality Models, Life Cycle, and Social Aspects. In ICT Innovations for Sustainability, Lorenz M. Hilty and Bernard Aebischer (Eds.). Advances in Intelligent Systems and Computing, Vol. 310. Springer, 191–205. https://doi.org/10.1007/978-3-319-09228-7_11
- Noureddine et al. (2012) Adel Noureddine, Aurelien Bourdon, Romain Rouvoy, and Lionel Seinturier. 2012. A preliminary study of the impact of software engineering on GreenIT. In First Int. Workshop on Green and Sustainable Software, GREENS 2012, Zurich, Switzerland, June 3, 2012, Rick Kazman et al. (Eds.). IEEE Computer Society, 21–27. https://doi.org/10.1109/GREENS.2012.6224251
- OECD (2009) OECD. 2009. Towards Green ICT Strategies: Assessing Policies and Programmes on ICT and the Environment. https://www.oecd-ilibrary.org/science-and-technology/towards-green-ict-strategies_222431651031.
- Ormazabal et al. (2014) Marta Ormazabal, Carmen Jaca, and Rogério Puga-Leal. 2014. Analysis and Comparison of Life Cycle Assessment and Carbon Footprint Software. In Proceedings of the Eighth International Conference on Management Science and Engineering Management, Jiuping Xu et al. (Eds.). Springer Berlin Heidelberg, Berlin, Heidelberg, 1521–1530.
- Pa et al. (2017) Noraini Che Pa, Faizal Karim, and Sa’adah Hassan. 2017. Dashboard system for measuring green software design. In Proceeding - 2017 3rd Int. Conf. on Science in Information Technology: Theory and Application of IT for Education, Industry and Society in Big Data Era, ICSITech 2017, Vol. 2018-January. 325 – 329. https://doi.org/10.1109/ICSITech.2017.8257133
- Pan and Rajan (2022) Rangeet Pan and Hridesh Rajan. 2022. Decomposing convolutional neural networks into reusable and replaceable modules. In Proceedings of the 44th Int. Conf. on Software Engineering (Pittsburgh, Pennsylvania) (ICSE ’22). Association for Computing Machinery, New York, NY, USA, 524–535. https://doi.org/10.1145/3510003.3510051
- Patterson et al. (2022) David A. Patterson, Joseph Gonzalez, Urs Hölzle, Quoc V. Le, Chen Liang, Lluis-Miquel Munguia, Daniel Rothchild, David R. So, Maud Texier, and Jeff Dean. 2022. The Carbon Footprint of Machine Learning Training Will Plateau, Then Shrink. Computer 55, 7 (2022), 18–28. https://doi.org/10.1109/MC.2022.3148714
- Penzenstadler et al. (2012) Birgit Penzenstadler, Veronika Bauer, Coral Calero, and Xavier Franch. 2012. Sustainability in software engineering: A systematic literature review. In 16th International Conference on Evaluation & Assessment in Software Engineering, EASE 2012, Ciudad Real, Spain, May 14-15, 2012. Proceedings, Maria Teresa Baldassarre et al. (Eds.). IET / IEEE Xplore, 32–41. https://doi.org/10.1049/IC.2012.0004
- Penzenstadler and Femmer (2013) Birgit Penzenstadler and Henning Femmer. 2013. A generic model for sustainability with process- and product-specific instances. In Proceedings of the 2013 Workshop on Green in/by Software Engineering, GIBSE ’13, Fukuoka, Japan, March 26, 2013, Somayeh Malakuti et al. (Eds.). ACM, 3–8. https://doi.org/10.1145/2451605.2451609
- Petersen et al. (2008) Kai Petersen, Robert Feldt, Shahid Mujtaba, and Michael Mattsson. 2008. Systematic mapping studies in software engineering. In Proceedings of the 12th International Conference on Evaluation and Assessment in Software Engineering (Italy) (EASE’08). BCS Learning & Development Ltd., Swindon, GBR, 68–77.
- Petersen et al. (2015) Kai Petersen, Sairam Vakkalanka, and Ludwik Kuzniarz. 2015. Guidelines for conducting systematic mapping studies in software engineering: An update. Information and Software Technology 64 (2015), 1–18. https://doi.org/10.1016/j.infsof.2015.03.007
- Ponsard et al. (2018) Christophe Ponsard, Renaud De Landtsheer, and Fabian Germeau. 2018. Building sustainable software for sustainable systems: case study of a shared pick-up and delivery service. In Proceedings of the 6th Int. Workshop on Green and Sustainable Software, GREENS@ICSE 2018, Gothenburg, Sweden, May 27, 2018, Ivano Malavolta et al. (Eds.). ACM, 31–34. https://doi.org/10.1145/3194078.3194083
- Popescu et al. (2022) Daniela Popescu, Mihai Dragomir, Sorin Popescu, and Diana Dragomir. 2022. Building Better Digital Twins for Production Systems by Incorporating Environmental Related Functions—Literature Analysis and Determining Alternatives. Applied Sciences 12, 17 (2022). https://doi.org/10.3390/app12178657
- Radersma (2022) Reinder Radersma. 2022. Green Coding: Reduce Your Carbon Footprint. ERCIM News 2022, 131 (2022). https://ercim-news.ercim.eu/en131/special/green-coding-reduce-your-carbon-footprint
- Raisian et al. (2021) Komeil Raisian, Jamaiah Yahaya, Siti Rohana Ahmad Ibrahim, Aziz Deraman, and Tumen Yunos. 2021. The Green Software Measurement Structure Based on Sustainability Perspective. In Int. Conf. on Electrical Engineering and Informatics, ICEEI 2021, Kuala Terengganu, Malaysia, October 12-13, 2021. IEEE, 1–6. https://doi.org/10.1109/ICEEI52609.2021.9611108
- Sachs et al. (2022) Jeffrey Sachs, Christian Kroll, Guillaume Lafortune, Grayson Fuller, and Finn Woelm. 2022. Sustainable development report 2022. Cambridge University Press.
- Salam and Khan (2016) Muhammad Salam and Siffat Ullah Khan. 2016. Developing green and sustainable software: Success factors for vendors. In 2016 7th IEEE International Conference on Software Engineering and Service Science (ICSESS). 1059–1062. https://doi.org/10.1109/ICSESS.2016.7883248
- Sambhanthan and Potdar (2016) Arunasalam Sambhanthan and Vidyasagar M. Potdar. 2016. Business sustainability frameworks: a survey. In 14th IEEE International Conference on Software Engineering Research, Management and Applications, SERA 2016, Towson, MD, USA, June 8-10, 2016, Yeong-Tae Song (Ed.). IEEE Computer Society, 171–177. https://doi.org/10.1109/SERA.2016.7516143
- Sharma et al. (2015) Sunil Kumar Sharma, P. K. Gupta, and Reza Malekian. 2015. Energy efficient software development life cycle - An approach towards smart computing. In 2015 IEEE Int. Conf. on Computer Graphics, Vision and Information Security (CGVIS). 1–5. https://doi.org/10.1109/CGVIS.2015.7449881
- Shenoy and Eeratta (2011) Sanath. S. Shenoy and Raghavendra Eeratta. 2011. Green software development model: An approach towards sustainable software development. In 2011 Annual IEEE India Conf. 1–6. https://doi.org/10.1109/INDCON.2011.6139638
- Siegmund et al. (2022) Norbert Siegmund, Johannes Dorn, Max Weber, Christian Kaltenecker, and Sven Apel. 2022. Green Configuration: Can Artificial Intelligence Help Reduce Energy Consumption of Configurable Software Systems? Computer 55, 3 (2022), 74–81. https://doi.org/10.1109/MC.2021.3120048
- Sikand et al. (2023) Samarth Sikand, Vibhu Saujanya Sharma, Vikrant Kaulgud, and Sanjay Podder. 2023. Green AI Quotient: Assessing Greenness of AI-based software and the way forward. In 38th IEEE/ACM Int. Conf. on Automated Software Engineering, ASE 2023. IEEE, 1828–1833. https://doi.org/10.1109/ASE56229.2023.00115
- Simon et al. (2023) Thibault Simon, Pierre Rust, Romain Rouvoy, and Joël Penhoat. 2023. Uncovering the Environmental Impact of Software Life Cycle. In Int. Conf. on ICT for Sustainability, ICT4S 2023, Rennes, France, June 5-9, 2023. IEEE, 176–187. https://doi.org/10.1109/ICT4S58814.2023.00026
- Sinha and Chandel (2014) Sunanda Sinha and S.S. Chandel. 2014. Review of software tools for hybrid renewable energy systems. Renewable and Sustainable Energy Reviews 32 (2014), 192–205. https://doi.org/10.1016/j.rser.2014.01.035
- Sudarmadi and Garniwa (2023) Didik Sudarmadi and Iwa Garniwa. 2023. A Review of Techno-economic Analysis and Size Optimization of Hybrid Renewable Energy Systems Using Homer. In 2023 4th International Conference on High Voltage Engineering and Power Systems (ICHVEPS). 401–406. https://doi.org/10.1109/ICHVEPS58902.2023.10257527
- Tee et al. (2014a) Mcxin Tee, Rusli Abdullah, and Salfarina Abdullah. 2014a. Toward Developing Green Software Development Model in Managing Knowledge of IT Practitioners for Sustaining Future Generation. In Knowledge Management Int. Conf. 2014 (KMICe).
- Tee et al. (2014b) Mcxin Tee, Rusli Abdullah, and Salfarina Abdullah. 2014b. Web-based knowledge management model for managing and sharing green knowledge of software development in community of practice. In 2014 8th. Malaysian Software Engineering Conf. (MySEC). 210–215. https://doi.org/10.1109/MySec.2014.6986016
- Venters et al. (2023) Colin C. Venters, Rafael Capilla, Elisa Yumi Nakagawa, Stefanie Betz, Birgit Penzenstadler, Tom Crick, and Ian Brooks. 2023. Sustainable software engineering: Reflections on advances in research and practice. Inf. Softw. Technol. 164 (2023), 107316. https://doi.org/10.1016/J.INFSOF.2023.107316
- Venters et al. (2018) Colin C. Venters, Rafael Capilla, Stefanie Betz, Birgit Penzenstadler, Tom Crick, Steve Crouch, Elisa Yumi Nakagawa, Christoph Becker, and Carlos Carrillo. 2018. Software sustainability: Research and practice from a software architecture viewpoint. J. Syst. Softw. 138 (2018), 174–188. https://doi.org/10.1016/J.JSS.2017.12.026
- Verdecchia et al. (2021) Roberto Verdecchia, Patricia Lago, Christof Ebert, and Carol de Vries. 2021. Green IT and Green Software. IEEE Softw. 38, 6 (2021), 7–15. https://doi.org/10.1109/MS.2021.3102254
- Verdecchia et al. (2023) Roberto Verdecchia, June Sallou, and Luis Cruz. 2023. A systematic review of Green AI. WIREs Data. Mining. Knowl. Discov. 13, 4 (2023). https://doi.org/10.1002/WIDM.1507
- Wedyan et al. (2023) Fadi Wedyan, Rachael Morrison, and Osama Sam Abuomar. 2023. Integration and Unit Testing of Software Energy Consumption. In Tenth Int. Conf. on Software Defined Systems, SDS 2023, San Antonio, TX, USA, October 23-25, 2023, Muhannad Quwaider et al. (Eds.). IEEE, 60–64. https://doi.org/10.1109/SDS59856.2023.10329262
- Wei et al. (2023) Xiaokai Wei, Sujan Kumar Gonugondla, Shiqi Wang, Wasi Uddin Ahmad, Baishakhi Ray, Haifeng Qian, Xiaopeng Li, Varun Kumar, Zijian Wang, Yuchen Tian, Qing Sun, Ben Athiwaratkun, Mingyue Shang, Murali Krishna Ramanathan, Parminder Bhatia, and Bing Xiang. 2023. Towards Greener Yet Powerful Code Generation via Quantization: An Empirical Study. In Proceedings of the 31st ACM Joint European Software Engineering Conf. and Symposium on the Foundations of Software Engineering, ESEC/FSE 2023, San Francisco, CA, USA, December 3-9, 2023, Satish Chandra et al. (Eds.). ACM, 224–236. https://doi.org/10.1145/3611643.3616302
- Wohlin (2014) Claes Wohlin. 2014. Guidelines for snowballing in systematic literature studies and a replication in software engineering. In 18th International Conference on Evaluation and Assessment in Software Engineering, EASE ’14, London, England, United Kingdom, May 13-14, 2014, Martin J. Shepperd et al. (Eds.). ACM, 38:1–38:10. https://doi.org/10.1145/2601248.2601268
- Wohlin et al. (2000) Claes Wohlin, Per Runeson, Martin Höst, Magnus C. Ohlsson, Bjöorn Regnell, and Anders Wesslén. 2000. Experimentation in software engineering: an introduction. Kluwer Academic Publishers, Norwell, MA, USA.
- Zhang et al. (2023) Peng Zhang, Yong Xiao, Yingyu Li, Xiaohu Ge, Guangming Shi, and Yang Yang. 2023. Toward net-zero carbon emissions in network AI for 6G and beyond. IEEE Communications Magazine 62, 4 (2023), 58–64. https://doi.org/10.1109/MCOM.003.2300175
- Zulfiqar et al. (2023) Muhammad Zulfiqar, Safdar Husain Tahir, Muhammad Rizwan Ullah, and Sadeen Ghafoor. 2023. Digitalized world and carbon footprints: does digitalization really matter for sustainable environment? Environ. Sci. Pollut. Res. 30 (2023), 88789–88802. https://doi.org/10.1007/s11356-023-28332-z